### v1

# Rigidity of Gradient Shrinking Ricci Solitons with Bach-like Flatness and Related Variational Formulas

Abstract
The classical Bach tensor in four dimensions can be expressed as a linear combination of two independent, symmetric, divergence-free, quadratic in curvature tensors U and V. Several classification results for gradient-shrinking Ricci solitons have been obtained under the assumption that the Bach tensor vanishes. We define a Bach-like tensor to be any other linear combination of U and V. We prove that, except along the Bach line, the vanishing of a Bach-like tensor forces a four-dimensional complete gradient-shrinking Ricci soliton to be either Einstein or isometric to the Gaussian soliton. Finally, we show that Bach-like tensors arise as Euler–Lagrange equations of a two-parameter family of quadratic curvature functionals and compute the corresponding first and second variation formulas.
1 Introduction and Main Results
A gradient Ricci soliton is a Riemannian manifold $(M^{n},g)$ together with a smooth potential function $f:M\to\mathbb{R}$ satisfying
$\mathrm{Rc}+\nabla^{2}f\;=\;\rho\,g,$ (1)
for $\rho$ a constant. The soliton is said to be shrinking if $\rho>0$, and it is common to scale $f$ so that $\rho=\frac{1}{2}$ for convenience. Such solitons arise as self-similar solutions of Hamilton’s Ricci flow, as models for type I singularities. Unless otherwise stated, our work will be in dimension $n=4$.
The classical Bach tensor was introduced by R. Bach [bach:1921] to study conformal relativity. It is defined in dimension 4 by:
$\mathrm{B}_{ij}=\nabla^{k}\nabla^{l}\mathrm{W}_{ijkl}+\frac{1}{2}\mathrm{R}^{kl}\mathrm{W}_{ikjl},$ (2)
where $\mathrm{W}_{ijkl}$ denotes the Weyl curvature and $R_{ij}$ the Ricci curvature. It is well-known that $\mathrm{B}_{ij}$ is conformally invariant, symmetric, trace-free, divergence-free, and quadratic in the Riemann curvature tensor. As seen in [bergman:2004] there are exactly two independent symmetric 22-tensors that are divergence-free and quadratic in curvature in four dimensions. These are the following:
$\mathrm{U}_{ij}=2\mathrm{R}_{ipjq}\mathrm{R}^{pq}+\Delta\mathrm{R}_{ij}-\frac{1}{2}|\mathrm{Rc}|^{2}g_{ij}-\mathrm{R}\,\mathrm{R}_{ij}-\frac{1}{2}\Delta\mathrm{R}\,g_{ij}+\frac{1}{4}\mathrm{R}^{2}\,g_{ij},$ (3)
$\mathrm{V}_{ij}=-\nabla_{i}\nabla_{j}\mathrm{R}+\Delta\mathrm{R}\,g_{ij}+\mathrm{R}\,\mathrm{R}_{ij}-\frac{1}{4}\mathrm{R}^{2}\,g_{ij}.$ (4)
One can express the Bach tensor as a combination of $U$ and $V$. In dimension 4 we have:
$\mathrm{B}_{ij}=\frac{1}{2}\mathrm{U}_{ij}+\frac{1}{6}\mathrm{V}_{ij}.$ (5)
We define a Bach-like tensor to be of the form:
$\mathfrak{B}_{ij}=\alpha\,\mathrm{U}_{ij}+\beta\,\mathrm{V}_{ij}$ (6)
for constants $(\alpha,\beta)\in\mathbb{R}^{2}\setminus\{\mathbf{0}\}$. If $\alpha/\beta=3$ then the Bach-like tensor is simply a rescaled Bach tensor. H.-D. Cao and Q. Chen showed in [caochen:2013] that if the Bach tensor vanishes on a complete gradient-shrinking Ricci soliton, then it must be rigid. Once we move beyond the Bach tensor itself, we can show that Bach-like flatness still implies rigidity, but without the index-heavy computations typical of the Bach-flat case. On the one hand, our results here are somewhat stronger in that only two classifications occur; on the other, this suggests that the Bach tensor behaves more like a genuine analytic obstruction, requiring more delicate analysis that ties directly into the geometry of the conformal Cotton tensor $\mathrm{D}_{ijk}$, which will be discussed later.
Theorem 1.1 (Vanishing Basis-Tensors).
Let $(M^{4},g,f)$ be a complete gradient-shrinking Ricci soliton satisfying (1). Then
  1. (i)
    $\mathrm{V}_{ij}\equiv 0$ if and only if $M^{4}$ is either Einstein or isometric to the Gaussian shrinking soliton;
  2. (ii)
    $\mathrm{U}_{ij}\equiv 0$ if and only if $M^{4}$ is either Einstein or isometric to the Gaussian shrinking soliton.
Theorem 1.2 (Bach-like Rigidity).
Let $(M^{4},g,f)$ be a complete gradient-shrinking Ricci soliton satisfying (1). Assume $(\alpha,\beta)\in\mathbb{R}^{2}\setminus\{\mathbf{0}\}$ such that $\beta\neq\alpha/3$. Then if
$\mathfrak{B}_{ij}=\alpha\,\mathrm{U}_{ij}+\beta\,\mathrm{V}_{ij}=0,$
$M^{4}$ is either Einstein or isometric to the Gaussian shrinking soliton. Conversely, every Einstein or Gaussian shrinking soliton satisfies $\mathfrak{B}_{ij}=0$. Hence, within the class of complete gradient-shrinking Ricci solitons, $\mathfrak{B}_{ij}=0$ is equivalent to $M^{4}$ being Einstein or isometric to the Gaussian soliton.
We will approach the proof of Theorem 1.2 by first considering the cone
$\mathcal{C}=\{(\alpha,\beta)\,\,|\,\,\alpha\geq 0\,\,\mathrm{and}\,\,\beta>\alpha/3\}\cup\{(\alpha,\beta)\,\,|\,\,\alpha\leq 0\,\,\mathrm{and}\,\,\beta<\alpha/3\},$
and primarily using weighted-integral identities from the $\mathrm{V}_{ij}$ tensor. For the region outside the cone we will take a different approach using more general weighted-integral identities. This result covers the entire $(\alpha,\beta)$ plane aside from the Bach-line and the origin.
2 Integral Identities
To proceed, we rely on a number of integral identities implied by $\Delta R=0$. The connection between harmonic scalar curvature and Bach-like tensors is rooted in the traces of U and V.
Lemma 2.1.
Let $U_{ij}$ and $V_{ij}$ be given by (3) and (4) respectively. Then
$\mathrm{tr}(\mathrm{U}_{ij})$ $=-\Delta R$
$\mathrm{tr}(\mathrm{V}_{ij})$ $=3\Delta R.$
So if $(\alpha,\beta)\in\mathcal{C}$, then $\mathfrak{B}_{ij}=0$ implies $\Delta R=0.$ We now derive the integral identities for both the unweighted and weighted volume elements.
Lemma 2.2 (Integral Identities).
Let $(M^{4},g,f)$ be a complete gradient-shrinking Ricci soliton satisfying (1) and assume that $\Delta\mathrm{R}=0$. Let $r$ be a regular value of $f$ and $\Omega_{r}=\{x\in M\,|\,f(x)\leq r\}$. Then:
  1. 1.)
    $\int_{\partial\Omega_{r}}\frac{1}{|\nabla f|}\langle\nabla\mathrm{R},\nabla f\rangle\,\mathrm{d}S=0,$
  2. 2.)
    $\int_{\Omega_{r}}\langle\nabla\mathrm{R},\nabla f\rangle\,\mathrm{d}V=0,$
  3. 3.)
    $\int_{\Omega_{r}}|\nabla\mathrm{R}|^{2}\,\mathrm{d}V=\int_{\partial\Omega_{r}}\frac{\mathrm{R}}{|\nabla f|}\langle\nabla\mathrm{R},\nabla f\rangle\,\mathrm{d}S=-\int_{\partial\Omega_{r}}|\nabla f|\langle\nabla\mathrm{R},\nabla f\rangle\,\mathrm{d}S,$
  4. 4.)
    $\int_{\Omega_{r}}\langle\nabla\mathrm{R},\nabla f\rangle\,\mathrm{e}^{-f}\,\mathrm{d}V=0,$
  5. 5.)
    $\int_{\Omega_{r}}\mathrm{R}\,\mathrm{Rc}(\nabla f,\nabla f)\,\mathrm{e}^{-f}\,\mathrm{d}V=\frac{1}{2}\int_{\partial\Omega_{r}}|\nabla f|\langle\nabla\mathrm{R},\nabla f\rangle\mathrm{e}^{-f}\,\mathrm{d}S+\frac{1}{2}\int_{\Omega_{r}}|\nabla\mathrm{R}|^{2}\,\mathrm{e}^{-f}\,\mathrm{d}V,$
  6. 6.)
    $\int_{\Omega_{r}}f\langle\nabla\mathrm{R},\nabla f\rangle\,\mathrm{e}^{-f}\,\mathrm{d}V=0.$
Proof.
First, we have:
$0=\int_{\Omega_{r}}\Delta R\,\mathrm{d}V=\int_{\partial\Omega_{r}}\frac{1}{|\nabla f|}\langle\nabla R,\nabla f\rangle\,\mathrm{d}S$
showing (1). Now we integrate by parts to get:
$\int_{\Omega_{r}}\langle\nabla\mathrm{R},\nabla f\rangle\,\mathrm{d}V$ $=\int_{\partial\Omega_{r}}\frac{f}{|\nabla f|}\langle\nabla\mathrm{R},\nabla f\rangle\,\mathrm{d}S-\int_{\Omega_{r}}f\Delta\mathrm{R}\,\mathrm{d}V$
$=r\int_{\partial\Omega_{r}}\frac{1}{|\nabla f|}\langle\nabla\mathrm{R},\nabla f\rangle\,\mathrm{d}S-\int_{\Omega_{r}}f\Delta\mathrm{R}\,\mathrm{d}V.$
As $\Delta R=0$ by assumption, and using (1) we see that both integrals must vanish, showing (2). To show (3) we integrate $R\Delta R$:
$0=\int_{\Omega_{r}}\mathrm{R}\Delta\mathrm{R}\,\mathrm{d}V=\int_{\partial\Omega_{r}}\frac{\mathrm{R}}{|\nabla f|}\langle\nabla\mathrm{R},\nabla f\rangle\,\mathrm{d}S-\int_{\Omega_{r}}|\nabla\mathrm{R}|^{2}\,\mathrm{d}V.$
And by using the identity $\mathrm{R}=f-|\nabla f|^{2}$ from normalizing $f$ as in [hamilton:1988], we get (3). Now we have:
$0$ $=\int_{\Omega_{r}}\Delta\mathrm{R}\,\mathrm{e}^{-f}\,\mathrm{d}V$
$=\int_{\partial\Omega_{r}}\frac{\mathrm{e}^{-f}}{|\nabla f|}\langle\nabla\mathrm{R},\nabla f\rangle\,\mathrm{d}S+\int_{\Omega_{r}}\langle\nabla\mathrm{R},\nabla f\rangle\,\mathrm{e}^{-f}\,\mathrm{d}V$
$=\mathrm{e}^{-r}\int_{\partial\Omega_{r}}\frac{1}{|\nabla f|}\langle\nabla\mathrm{R},\nabla f\rangle\,\mathrm{d}S+\int_{\Omega_{r}}\langle\nabla\mathrm{R},\nabla f\rangle\,\mathrm{e}^{-f}\,\mathrm{d}V$
$=\int_{\Omega_{r}}\langle\nabla\mathrm{R},\nabla f\rangle\,\mathrm{e}^{-f}\,\mathrm{d}V$
showing (4). To show (5) we integrate $R\Delta R$ with the weighted volume $\mathrm{e}^{-f}\,\mathrm{d}V$:
$0$ $=\int_{\Omega_{r}}\mathrm{R}\Delta\mathrm{R}\,\mathrm{e}^{-f}\,\mathrm{d}V$
$=\int_{\partial\Omega_{r}}\frac{\mathrm{R}}{|\nabla f|}\langle\nabla\mathrm{R},\nabla f\rangle\,\mathrm{e}^{-f}\,\mathrm{d}S-\int_{\Omega_{r}}\langle\nabla\big(\mathrm{R}\,\mathrm{e}^{-f}\big),\nabla\mathrm{R}\rangle\,\mathrm{d}V$
$=\int_{\partial\Omega_{r}}\frac{\mathrm{R}}{|\nabla f|}\langle\nabla\mathrm{R},\nabla f\rangle\,\mathrm{e}^{-f}\,\mathrm{d}S-\int_{\Omega_{r}}\Big(|\nabla\mathrm{R}|^{2}-R\langle\nabla\mathrm{R},\nabla f\rangle\Big)\,\mathrm{e}^{-f}\,\mathrm{d}V$
$=-\int_{\partial\Omega_{r}}|\nabla f|\langle\nabla\mathrm{R},\nabla f\rangle\,\mathrm{e}^{-f}\,\mathrm{d}S-\int_{\Omega_{r}}\Big(|\nabla\mathrm{R}|^{2}-2\mathrm{R}\,\mathrm{Rc}(\nabla f,\nabla f)\Big)\,\mathrm{e}^{-f}\,\mathrm{d}V.$
And lastly for (6), we have:
$\int_{\Omega_{r}}f\langle\nabla\mathrm{R},\nabla f\rangle\,\mathrm{e}^{-f}\,\mathrm{d}V$ $=-\int_{\Omega_{r}}\langle f\nabla\,\mathrm{R},\nabla\mathrm{e}^{-f}\rangle\,\mathrm{d}V$
$=-\int_{\partial\Omega_{r}}\frac{f}{|\nabla f|}\langle\nabla\mathrm{R},\nabla f\rangle\,\mathrm{e}^{-f}\,\mathrm{d}S+\int_{\Omega_{r}}\mathrm{div}(f\,\nabla\mathrm{R})\mathrm{e}^{-f}\,\mathrm{d}V$
$=-r\mathrm{e}^{-r}\int_{\partial\Omega_{r}}\frac{1}{|\nabla f|}\langle\nabla\mathrm{R},\nabla f\rangle\,\mathrm{d}S+\int_{\Omega_{r}}\mathrm{div}(f\,\nabla\mathrm{R})\mathrm{e}^{-f}\,\mathrm{d}V$
$=\int_{\Omega_{r}}\Big(\langle\nabla\mathrm{R},\nabla f\rangle+f\,\Delta\mathrm{R}\Big)\mathrm{e}^{-f}\,\mathrm{d}V$
$=0.$
3 The Vanishing V-tensor
The tensor $\mathrm{V}_{ij}$ has fewer and less complicated terms than $\mathrm{U}_{ij}$. We will obtain integral identities by assuming $\mathrm{V}_{ij}=0$, and use these to prove Theorem 1.1. We first consider the Hessian term:
Lemma 3.1.
Let $(M^{4},g,f)$ be a complete gradient-shrinking Ricci soliton satisfying (1) and assume that $\Delta R=0$. Then if $r$ is a regular value of $f$ and $\Omega_{r}=\{x\in M\,|\,f(x)\leq r\}$, we have:
$\int_{\Omega_{r}}\nabla^{2}\mathrm{R}(\nabla f,\nabla f)\,\mathrm{e}^{-f}\,\mathrm{d}V=-\mathrm{e}^{-r}\int_{\Omega_{r}}|\nabla\mathrm{R}|^{2}\,\mathrm{d}V+\frac{1}{2}\int_{\Omega_{r}}|\nabla\mathrm{R}|^{2}\,\mathrm{e}^{-f}\,\mathrm{d}V.$
Proof.
We integrate:
$\int_{\Omega_{r}}\nabla^{2}$ $\mathrm{R}(\nabla f,\nabla f)\,\mathrm{e}^{-f}\,\mathrm{d}V$
$=\int_{\Omega_{r}}\langle\nabla\langle\nabla\mathrm{R},\nabla f\rangle,\nabla f\rangle\,\mathrm{e}^{-f}\,\mathrm{d}V-\frac{1}{2}\int_{\Omega_{r}}\langle\nabla\mathrm{R},\nabla f\rangle\,\mathrm{e}^{-f}\,\mathrm{d}V$
$\hskip 36.135pt+\frac{1}{2}\int_{\Omega_{r}}|\nabla\mathrm{R}|^{2}\,\mathrm{e}^{-f}\,\mathrm{d}V$
$=-\int_{\Omega_{r}}\langle\nabla\langle\nabla\mathrm{R},\nabla f\rangle,\nabla\mathrm{e}^{-f}\rangle\,\mathrm{d}V+\frac{1}{2}\int_{\Omega_{r}}|\nabla\mathrm{R}|^{2}\,\mathrm{e}^{-f}\,\mathrm{d}V$
$=\int_{\partial\Omega_{r}}|\nabla f|\langle\nabla\mathrm{R},\nabla f\rangle\mathrm{e}^{-f}\,\mathrm{d}S+\int_{\Omega_{r}}\Delta(\mathrm{e}^{-f})\langle\nabla\mathrm{R},\nabla f\rangle\,\mathrm{d}V$
$\hskip 36.135pt+\frac{1}{2}\int_{\Omega_{r}}|\nabla\mathrm{R}|^{2}\,\mathrm{e}^{-f}\,\mathrm{d}V$
$=\int_{\partial\Omega_{r}}|\nabla f|\langle\nabla\mathrm{R},\nabla f\rangle\mathrm{e}^{-f}\,\mathrm{d}S+\int_{\Omega_{r}}(|\nabla f|^{2}-\Delta f)\langle\nabla\mathrm{R},\nabla f\rangle\mathrm{e}^{-f}\,\mathrm{d}V$
$\hskip 36.135pt+\frac{1}{2}\int_{\Omega_{r}}|\nabla\mathrm{R}|^{2}\,\mathrm{e}^{-f}\,\mathrm{d}V$
$=\int_{\partial\Omega_{r}}|\nabla f|\langle\nabla\mathrm{R},\nabla f\rangle\mathrm{e}^{-f}\,\mathrm{d}S+\int_{\Omega_{r}}(f-2)\langle\nabla\mathrm{R},\nabla f\rangle\mathrm{e}^{-f}\,\mathrm{d}V$
$\hskip 36.135pt+\frac{1}{2}\int_{\Omega_{r}}|\nabla\mathrm{R}|^{2}\,\mathrm{e}^{-f}\,\mathrm{d}V.$
Lemma 7.1 (1) and (3) shows that the integral of $(f-2)\langle\nabla\mathrm{R},\nabla f\rangle$ vanishes, giving us our desired result. ∎
It will be necessary to control the $|\nabla\mathrm{R}|^{2}$ integral as we allow $r\to\infty$. Work already exists for controlling the integral of $|\mathrm{Rc}|^{2}$, which we are able to translate to our problem.
Lemma 3.2.
Let $(M^{n},g_{ij},f)$ be a complete gradient-shrinking Ricci soliton satisfying (1). Then:
$\int_{M}|\nabla\mathrm{R}|^{2}\mathrm{e}^{-\alpha f}\,\mathrm{d}V<\infty$
for any $\alpha>0$.
Proof.
By the Cauchy-Schwarz inequality we have:
$\frac{1}{4}|\nabla\mathrm{R}|^{2}=|\mathrm{Rc}(\nabla f)|^{2}\leq|\mathrm{Rc}|^{2}|\nabla f|^{2}$
By H.-D. Cao and D. Zhou’s growth estimates for the potential function [caozhou:2010], we have:
$|\nabla f|^{2}(x)\leq\frac{1}{4}(d(x)+c_{2})^{2}$
[…]
$\int_{M}|\mathrm{Rc}|^{2}|\nabla f|^{2}\mathrm{e}^{-\alpha f}\,\mathrm{d}V=\int_{K}|\mathrm{Rc}|^{2}|\nabla f|^{2}\mathrm{e}^{-\alpha f}\,\mathrm{d}V+\int_{M\backslash K}|\mathrm{Rc}|^{2}|\nabla f|^{2}\mathrm{e}^{-\alpha f}\,\mathrm{d}V.$
Clearly the first integral over the compact set $K$ is finite. Now we have for the second integral:
$\begin{split}\int_{M\backslash K}|\mathrm{Rc}|^{2}|\nabla f|^{2}\mathrm{e}^{-\alpha f}\,\mathrm{d}V&=\int_{M\backslash K}(|\mathrm{Rc}|^{2}\mathrm{e}^{-\alpha f/2})|\nabla f|^{2}\mathrm{e}^{-\alpha f/2}\,\mathrm{d}V\\ &\leq\int_{M\backslash K}|\mathrm{Rc}|^{2}\mathrm{e}^{-\alpha f/2}\,\mathrm{d}V.\end{split}$
By O. Munteanu and N. Sesum’s results in [munteanusesum:2013], we have:
$\int_{M}|\mathrm{Rc}|^{2}\mathrm{e}^{-\lambda f}\,\mathrm{d}V<\infty,$
for any $\lambda>0$, so after setting $\lambda=\frac{\alpha}{2}$ the result follows. ∎
Lemma 3.3.
Let $(M^{n},g_{ij},f)$ be a complete gradient-shrinking Ricci soliton satisfying (1). Then if $\Omega_{r}=\{f\leq r\}$ for regular values $r>0$ of $f$ and $\alpha>0$:
$\lim_{r\rightarrow\infty}\Big(\mathrm{e}^{-\alpha r}\int_{\Omega_{r}}|\nabla\mathrm{R}|^{2}\,\mathrm{d}V\Big)=0.$
Proof.
We first apply Lemma 3.2:
$\int_{M}|\nabla\mathrm{R}|^{2}\mathrm{e}^{-\alpha f/2}\,\mathrm{d}V<\infty$
Since $f\leq r$ on $\Omega_{r}$, we also have $\mathrm{e}^{-\alpha r/2}\leq\mathrm{e}^{-\alpha f/2}$:
$\frac{\mathrm{e}^{-\alpha r}}{2}\int_{\Omega_{r}}|\nabla\mathrm{R}|^{2}\,\mathrm{d}V=\frac{\mathrm{e}^{-\alpha r/2}}{2}\int_{\Omega_{r}}|\nabla\mathrm{R}|^{2}\mathrm{e}^{-\alpha r/2}\,\mathrm{d}V\leq\frac{\mathrm{e}^{-\alpha r/2}}{2}\int_{\Omega_{r}}|\nabla\mathrm{R}|^{2}\mathrm{e}^{-\alpha f/2}\,\mathrm{d}V.$
Now let $r\rightarrow\infty$, giving us the desired result:
$\lim_{r\rightarrow\infty}\Big(\frac{\mathrm{e}^{-\alpha r}}{2}\int_{\Omega_{r}}|\nabla\mathrm{R}|^{2}\,\mathrm{d}V\Big)\leq\lim_{r\rightarrow\infty}\Big(\frac{\mathrm{e}^{-\alpha r/2}}{2}\int_{\Omega_{r}}|\nabla\mathrm{R}|^{2}\mathrm{e}^{-\alpha f/2}\,\mathrm{d}V\Big)=0.$
4 Proof of Theorem 1.1(i)
Proof.
Let $\Omega_{r}=\{x\in M\,|\,f(x)\leq r\}$ where $r$ is a regular value of $f$. By Lemma 2.1, $\mathrm{V}\equiv 0$ implies $\Delta\mathrm{R}=0$. By Lemma 2.2(5) and (3) we have:
$\int_{\Omega_{r}}\mathrm{R}\,\mathrm{Rc}(\nabla f,\nabla f)\,\mathrm{e}^{-f}\,\mathrm{d}V$ $=-\frac{1}{2}\int_{\Omega_{r}}\mathrm{R}\Delta_{f}\mathrm{R}\mathrm{e}^{-f}\,\mathrm{d}V$
$=-\frac{1}{2}\int_{\Omega_{r}}|\nabla\mathrm{R}|^{2}\,\mathrm{e}^{-r}\,\mathrm{d}V+\frac{1}{2}\int_{\Omega_{r}}|\nabla\mathrm{R}|^{2}\,\mathrm{e}^{-f}\,\mathrm{d}V.$
And using Lemma 3.1, we compute the integral of $V(\nabla f,\nabla f)$ with respect to the weighted volume element $\mathrm{e}^{-f}\,\mathrm{d}V$, and we note that the second term involving $\Delta R$ again vanishes:
$\int_{\Omega_{r}}\mathrm{V}($ $\nabla f,\nabla f)\,\mathrm{e}^{-f}\,\mathrm{d}V$
$=-\int_{\Omega_{r}}\Big(\nabla^{2}\mathrm{R}(\nabla f,\nabla f)-\mathrm{R}\,\mathrm{Rc}(\nabla f,\nabla f)+\frac{1}{4}\mathrm{R}^{2}|\nabla f|^{2}\Big)\,\mathrm{e}^{-f}\,\mathrm{d}V$
$=\frac{\mathrm{e}^{-r}}{2}\int_{\Omega_{r}}|\nabla\mathrm{R}|^{2}\,\mathrm{d}V-\frac{1}{4}\int_{\Omega_{r}}\mathrm{R}^{2}|\nabla f|^{2}\,\mathrm{e}^{-f}\,\mathrm{d}V.$
Letting $r\rightarrow\infty$, by Lemma 3.3 the first integral vanishes. We obtain:
$\int_{M}\mathrm{V}(\nabla f,\nabla f)\,\mathrm{e}^{-f}\,\mathrm{d}V=-\int_{M}\frac{1}{4}\mathrm{R}^{2}|\nabla f|^{2}\,\mathrm{e}^{-f}\,\mathrm{d}V.$ (7)
Assuming $\mathrm{V}_{ij}=0$, for all $p\in M$ we must have either $\mathrm{R}=0$ or $|\nabla f|=0$. By the work of S. Pigola, M. Rimoldi, and A. Setti [pigolarimoldi:2011], either $R>0$ everywhere, hence $\nabla f=0$ and the manifold is Einstein, or if $R=0$ anywhere, then the soliton must be isometric to the Gaussian soliton on $\mathbb{R}^{4}$.
For the converse, suppose $(M^{4},g,f)$ is Einstein. Then as the scalar curvature is constant, we have:
$\mathrm{V}_{ij}$ $=\frac{1}{4}\mathrm{R}^{2}g_{ij}-\mathrm{R}\,\mathrm{R}_{ij}$
$=-\mathrm{R}(\mathrm{R}_{ij}-\frac{1}{4}\mathrm{R}\,g_{ij})$
$=0.$
If the manifold is Gaussian, then $\mathrm{R}=0$ and again $\mathrm{V}=0$. ∎
5 Weighted Bach Integral
In [caochen:2013] H.-D. Cao and Q. Chen consider how the Bach tensor links to the Cotton tensor of a conformal metric, called $\mathrm{D}_{ijk}$. The norm of $\mathrm{D}_{ijk}$ contains rich geometric information, and they were able to use it to prove their classification results. We will link Bach-like tensors to $\mathrm{D}_{ijk}$ as well. Recall that the Cotton tensor is given by:
$\mathrm{C}_{ijk}=-\frac{n-2}{n-3}\nabla_{l}\mathrm{W}_{ijkl},$
where $\mathrm{W}_{ijkl}$ is the Weyl curvature. The Cotton tensor is skew-symmetric in the first two indices and trace-free in any two indices. That is,
$\mathrm{C}_{ijk}=-\mathrm{C}_{jik}\quad\mathrm{and}\quad g^{ij}\,\mathrm{C}_{ijk}=g^{ik}\,\mathrm{C}_{ijk}=0.$
The Cotton tensor for the conformal metric $\widehat{g}_{ij}=\mathrm{e}^{\frac{1}{2}f}\,g_{ij}$ is given by:
$\mathrm{D}_{ijk}=\mathrm{C}_{ijk}+\mathrm{W}_{ijkl}\,\nabla_{l}f.$
On a complete gradient-shrinking Ricci soliton, $\mathrm{D}_{ijk}$ can be expressed as:
$D_{ijk}$ $=\frac{1}{n-2}\big(\mathrm{R}_{jk}\nabla_{i}f-\mathrm{R}_{ik}\nabla_{j}f\big)+\frac{1}{2(n-1)(n-2)}\big(g_{jk}\,\nabla_{i}\mathrm{R}-g_{ik}\,\nabla_{j}\mathrm{R}\big)$
$\quad-\frac{1}{(n-1)(n-2)}\mathrm{R}\,\big(g_{jk}\,\nabla_{i}f-g_{ik}\,\nabla_{j}f\big).$
H.-D. Cao and Q. Chen [caochen:2013] showed how the Bach tensor can be expressed in terms of $\mathrm{C}_{ijk}$ and $\mathrm{D}_{ijk}$:
$\mathrm{B}_{ij}=-\frac{1}{n-2}\Big(\nabla_{k}\mathrm{D}_{ikj}+\frac{n-3}{n-2}\mathrm{C}_{jli}\nabla_{l}f\Big).$
We can now link together the integral of the Bach tensor contracted with the gradient of the potential function to the norm of $\mathrm{D}_{ijk}$.
Lemma 5.1.
Let $(M^{n},g_{ij},f)$ be a complete gradient-shrinking Ricci soliton satisfying (1). Then
$\int_{M}\mathrm{B}(\nabla f,\nabla f)e^{-f}\,\mathrm{d}V=-\frac{1}{2}\int_{M}|\mathrm{D}_{ijk}|^{2}e^{-f}\,\mathrm{d}V.$
Proof.
Let $r$ be a regular value of $f$ and $\Omega_{r}=\Omega_{r}=\{f\leq r\}$. We use the antisymmetric properties of the Cotton and $\mathrm{D}_{ijk}$ tensor:
$(n-2)\int_{\Omega_{r}}$ $\mathrm{B}_{ij}\nabla_{i}f\nabla_{j}f\,\mathrm{e}^{-f}\,\mathrm{d}V$
$=-\int_{\Omega_{r}}\nabla_{k}\mathrm{D}_{ikj}\nabla_{i}f\nabla_{j}f\mathrm{e}^{-f}\,\mathrm{d}V-\frac{n-3}{n-2}\int_{\Omega_{r}}\mathrm{C}_{jki}\nabla_{i}f\nabla_{j}f\nabla_{k}f\,\mathrm{e}^{-f}\,\mathrm{d}V$
$=-\int_{\partial\Omega_{r}}\mathrm{D}_{ikj}\nabla_{i}f\nabla_{j}f\nabla_{k}f\cdot|\nabla f|^{-1}\,\mathrm{e}^{-f}\,\mathrm{d}S$
$\quad+\int_{\Omega_{r}}\mathrm{D}_{ikj}(\mathrm{R}_{ik}\nabla_{j}f+\mathrm{R}_{kj}\nabla_{i}f)\,\mathrm{e}^{-f}\,\mathrm{d}V$
$\quad-\int_{M}\mathrm{D}_{ikj}\nabla_{i}f\nabla_{j}f\nabla_{k}f\,\mathrm{e}^{-f}\,\mathrm{d}V$
$=\int_{\Omega_{r}}\mathrm{D}_{ikj}(\mathrm{R}_{ik}\nabla_{j}f+\mathrm{R}_{kj}\nabla_{i}f)\,\mathrm{e}^{-f}\,\mathrm{d}V$
$=-\frac{1}{2}\int_{\Omega_{r}}\mathrm{D}_{ijk}(\mathrm{R}_{ik}\nabla_{j}f-\mathrm{R}_{kj}\nabla_{i}f)\,\mathrm{e}^{-f}\,\mathrm{d}V$
$=-\frac{(n-2)}{2}\int_{\Omega_{r}}|\mathrm{D}_{ijk}|^{2}\,\mathrm{e}^{-f}\,\mathrm{d}V.$
6 Proof of Theorem 1.2
First observe that for any $(\alpha,\beta)\in\mathcal{C}$, and a Bach-like tensor $\mathfrak{B}_{ij}=\alpha\,\mathrm{U}_{ij}+\beta\,\mathrm{V}_{ij}$ we can write:
$\mathfrak{B}_{ij}=2\alpha\,\mathrm{B}_{ij}+\Big(\beta-\frac{\alpha}{3}\Big)\,\mathrm{V}_{ij}.$
Assuming $\mathfrak{B}_{ij}=0$ and taking a trace we have:
$g^{ij}\,\mathfrak{B}_{ij}=(3\beta-\alpha)\Delta\mathrm{R}=0\implies\Delta\mathrm{R}=0,$
by Lemma 2.1. Then we have:
$\int_{M}\mathfrak{B}(\nabla f,\nabla f)\,\mathrm{e}^{-f}\,\mathrm{d}V$ $=-\alpha\int_{M}|\mathrm{D}_{ijk}|^{2}\,\mathrm{e}^{-f}\,\mathrm{d}V-\frac{1}{4}\Big(\beta-\frac{\alpha}{3}\Big)\int_{M}\mathrm{R}^{2}\,|\nabla f|^{2}\,\mathrm{e}^{-f}\,\,\mathrm{d}V=0.$
Since $(\alpha,\beta)\in\mathcal{C}$, we have two cases. If $\alpha>0$ and $\beta>\alpha/3$, then we must have $|\mathrm{D}_{ijk}|^{2}(p)=0$ and $\mathrm{R}^{2}(p)|\nabla f(p)|^{2}=0$ for all $p\in M$. Then we can conclude that $M$ is either Einstein or isometric to the Gaussian soliton as in the proof of Theorem 1.1. The argument for when $\alpha<0$ and $\beta<\alpha/3$ is analogous.
7 The Vanishing U-tensor
In order to extend the $(\alpha,\beta)$-cone in Theorem 1.2, we need to consider the implications of $\mathrm{U}_{ij}=0$. We attempt the same style of argument as we did with $\mathrm{V}_{ij}=0$. Define a weighted volume element of the form
$\mathrm{d}V_{c}(\mathbf{x})=\mathrm{e}^{-cf(\textbf{x})}\,\mathrm{d}V\quad\mathrm{where}\quad c>0.$
Lemma 7.1.
Let $(M^{4},f,g)$ be a complete gradient-shrinking Ricci soliton satisfying (1), normalized so that $\mathrm{R}+|\nabla f|^{2}=f$. Let $r$ be a regular value of $f$ and define the sub-level set $\Omega_{r}=\{x\in M\,|\,f(x)\leq r\}$. If $c>0$ is a constant and $\Delta\mathrm{R}=0$, then we have the following:
  1. (1.)
    $\int_{\Omega_{r}}\langle\nabla\mathrm{R},\nabla f\rangle\,\mathrm{d}V_{c}=0$,
  2. (2.)
    $\int_{\Omega_{r}}\mathrm{R}\langle\nabla\mathrm{R},\nabla f\rangle\,\mathrm{d}V_{c}=\frac{1}{c}\int_{\Omega_{r}}|\nabla\mathrm{R}|^{2}\,\mathrm{d}V_{c}-\frac{\mathrm{e}^{-c\,r}}{c}\int_{\Omega_{r}}|\nabla\mathrm{R}|^{2}\,\mathrm{d}V$,
  3. (3.)
    $\int_{\Omega_{r}}f\,\langle\nabla R,\nabla f\rangle\,\mathrm{d}V_{c}=0$.
Proof.
First we have:
$0$ $=\int_{\Omega_{r}}\Delta\mathrm{R}\mathrm{e}^{-cf}\,\mathrm{d}V$
$=\int_{\partial\Omega_{r}}\frac{1}{|\nabla f|}\langle\nabla\mathrm{R},\nabla f\rangle\mathrm{e}^{-cf}\,\mathrm{d}S+c\int_{\Omega_{r}}\langle\nabla\mathrm{R},\nabla f\rangle\mathrm{e}^{-cf}\,\,\mathrm{d}V$
$=c\int_{\Omega_{r}}\langle\nabla\mathrm{R},\nabla f\rangle\mathrm{e}^{-f}\,\,\mathrm{d}V.$
Next, we have:
$0$ $=\int_{\Omega_{r}}\mathrm{R}\Delta\mathrm{R}\,\mathrm{e}^{-cf}\,\mathrm{d}V$
$=\int_{\partial\Omega_{r}}\frac{\mathrm{R}}{|\nabla f|}\langle\nabla\mathrm{R},\nabla f\rangle\,\mathrm{e}^{-cf}\,\mathrm{d}S-\int_{\Omega_{r}}\Big(|\nabla\mathrm{R}|^{2}-c\,\mathrm{R}\,\langle\nabla\mathrm{R},\nabla f\rangle\Big)\mathrm{e}^{-cf}\,\mathrm{d}V.$
Hence,
$\int_{\Omega_{r}}\mathrm{R}\,\langle\nabla\mathrm{R},\nabla f\rangle\,\mathrm{e}^{-cf}\,\mathrm{d}V=\frac{1}{c}\int_{\Omega_{r}}|\nabla\mathrm{R}|^{2}\mathrm{e}^{-cf}\,\mathrm{d}V-\frac{\mathrm{e}^{-c\,r}}{c}\int_{\Omega_{r}}|\nabla\mathrm{R}|^{2}\,\mathrm{d}V,$
where we used 2.2(3). And the last critical identity is:
$\int_{\Omega_{r}}f$ $\langle\nabla\mathrm{R},\nabla f\rangle\mathrm{e}^{-cf}\,\mathrm{d}V$
$=-\frac{1}{c}\int_{\Omega_{r}}\langle f\,\nabla\mathrm{R},\nabla(\mathrm{e}^{-cf})\rangle\,\mathrm{d}V$
$=-\frac{1}{c}\Big[\int_{\partial\Omega_{r}}\frac{f}{|\nabla f|}\langle\nabla\mathrm{R},\nabla f\rangle\mathrm{e}^{-cf}\,\mathrm{d}S-\int_{\Omega_{r}}\mathrm{div}(f\nabla\mathrm{R})\mathrm{e}^{-cf}\,\mathrm{d}V\Big]$
$=-\frac{1}{c}\Big[r\,\mathrm{e}^{-cr}\int_{\partial\Omega_{r}}\frac{1}{|\nabla f|}\langle\nabla\mathrm{R},\nabla f\rangle\mathrm{e}^{-cf}\,\mathrm{d}S-\int_{\Omega_{r}}\mathrm{div}(f\,\nabla\mathrm{R})\mathrm{e}^{-cf}\,\mathrm{d}V\Big]$
$=\frac{1}{c}\int_{\Omega_{r}}\Big(\langle\nabla\mathrm{R},\nabla f\rangle+f\Delta\mathrm{R}\Big)\mathrm{e}^{-cf}\,\mathrm{d}V$
$=0.$
We first state a quick consequence of Bochner’s formula, and then we will proceed with the main integral identity.
Lemma 7.2.
Let $(M^{4},f,g)$ be a complete gradient-shrinking Ricci soliton satisfying (1), normalized so that $\mathrm{R}+|\nabla f|^{2}=f$. Then if $\Delta\mathrm{R}=0$:
$2|\mathrm{Rc}|^{2}=\mathrm{R}+\langle\nabla\mathrm{R},\nabla f\rangle$
Proof.
By Bochner’s formula, we have:
$\frac{1}{2}\Delta|\nabla f|^{2}$ $=\langle\nabla(\Delta f),\nabla f\rangle+|\mbox{Hess}(f)|^{2}+\mathrm{Rc}(\nabla f,\nabla f)$
$\frac{1}{2}\Delta(f-\mathrm{R})$ $=\langle\nabla(2-\mathrm{R}),\nabla f\rangle+|\mbox{Hess}(f)|^{2}+\frac{1}{2}\langle\nabla\mathrm{R},\nabla f\rangle$
$\frac{1}{2}(2-\mathrm{R})$ $=-\frac{1}{2}\langle\nabla\mathrm{R},\nabla f\rangle+|\frac{1}{2}g-\mathrm{Rc}|^{2}$
$1-\frac{\mathrm{R}}{2}$ $=-\frac{1}{2}\langle\nabla\mathrm{R},\nabla f\rangle+1-\mathrm{R}+|\mathrm{Rc}|^{2}$
$\frac{\mathrm{R}}{2}$ $=-\frac{1}{2}\langle\nabla\mathrm{R},\nabla f\rangle+|\mathrm{Rc}|^{2}.$
From here our claim clearly follows. ∎
Lemma 7.3.
Let $(M^{4},g,f)$ be a complete gradient-shrinking Ricci soliton satisfying (1). Assume that $\Delta\mathrm{R}\equiv 0$, $r>0$ is a regular value of $f$, and $\Omega_{r}=\{f\leq r\}$. Then
$\int_{\Omega_{r}}\mathrm{U}(\nabla f,\nabla f)\,\mathrm{d}V_{c}=\frac{1}{4}\int_{\Omega_{r}}\big(\mathrm{R}^{2}-2\,|\mathrm{Rc}|^{2}\big)\,|\nabla f|^{2}\,\mathrm{d}V_{c}.$
Moreover, this can be rewritten as
$\int_{\Omega_{r}}\mathrm{U}(\nabla f,\nabla f)\,\mathrm{d}V_{c}=\frac{1}{4}\int_{\Omega_{r}}\big(\mathrm{R}^{2}\,|\nabla f|^{2}-\mathrm{R}\,|\nabla f|^{2}+\frac{1}{c}|\nabla\mathrm{R}|^{2}\big)\,\mathrm{d}V_{c}-\frac{\mathrm{e}^{-cr}}{4c}\int_{\Omega_{r}}|\nabla\mathrm{R}|^{2}\,\mathrm{d}V.$
Proof.
One can easily check that on a soliton we have
$U_{ij}=\mathrm{R}_{ij}+\nabla_{k}\mathrm{R}_{ij}\nabla_{k}f-\frac{1}{2}|\mathrm{Rc}|^{2}g_{ij}-\mathrm{R}\,\mathrm{R}_{ij}-\frac{1}{2}\Delta\mathrm{R}g_{ij}+\frac{\mathrm{R}^{2}}{4}g_{ij},$
as well as the soliton identities
  1. (a)
    $\mathrm{\mathrm{Rc}}(\nabla f)=\frac{1}{2}\nabla\mathrm{R}\implies\mathrm{\mathrm{Rc}}(\nabla f,\nabla f)=\frac{1}{2}\langle\nabla\mathrm{R},\nabla f\rangle$,
  2. (b)
    $\Delta f=2-\mathrm{R}$,
  3. (c)
    $\nabla^{2}f(\nabla f)=\frac{1}{2}\nabla f-\frac{1}{2}\nabla\mathrm{R}$.
The second term in our expression for $U$ is simplified by looking at the following divergence:
$\nabla_{k}(\mathrm{R}_{ij}$ $\nabla_{i}f\nabla_{j}f\nabla_{k}f\,\mathrm{e}^{-cf})$
$=\nabla_{k}\mathrm{R}_{ij}\nabla_{i}f\nabla_{j}f\nabla_{k}f\mathrm{e}^{-cf}+2\,\mathrm{R}_{ij}\nabla_{i}f\nabla_{j}\nabla_{k}f(\nabla_{k}f)\mathrm{e}^{-cf}$
$\quad+\Delta f\,\mathrm{R}_{ij}\nabla_{i}f\nabla_{j}f\mathrm{e}^{-cf}-c|\nabla f|^{2}\mathrm{R}_{ij}\nabla_{i}f\nabla_{j}f\,\mathrm{e}^{-cf}.$
Now by the divergence theorem we have
$\int_{\Omega_{r}}$ $\nabla_{k}\mathrm{R}_{ij}\nabla_{i}f\nabla_{j}f\nabla_{k}f\,\mathrm{e}^{-cf}\,\mathrm{d}V$
$=\int_{\Omega_{r}}\mathrm{div}(\frac{1}{2}\langle\nabla\mathrm{R},\nabla f\rangle\nabla_{k}f\,\mathrm{e}^{-cf})\,\mathrm{d}V+\frac{1}{2}\int_{\Omega_{r}}\mathrm{R}\langle\nabla\mathrm{R},\nabla f\rangle\,\mathrm{e}^{-cf}\,\mathrm{d}V$
$\quad-\frac{1}{2}\int_{\Omega_{r}}|\nabla\mathrm{R}|^{2}\,\mathrm{e}^{-cf}\,\mathrm{d}V+\frac{\mathrm{e}^{-cr}}{2}\int_{\Omega_{r}}|\nabla\mathrm{R}|^{2}\,\mathrm{d}V$
$\quad-\int_{\Omega_{r}}\langle\nabla\mathrm{R},\frac{1}{2}\nabla f-\frac{1}{2}\nabla\mathrm{R}\rangle\,\mathrm{e}^{-cf}\,\mathrm{d}V$
$=\frac{1}{2}\int_{\partial\Omega_{r}}|\nabla f|\langle\nabla\mathrm{R},\nabla f\rangle\,\mathrm{e}^{-cf}\,\mathrm{d}S+\frac{1}{2c}\int_{\Omega_{r}}|\nabla\mathrm{R}|^{2}\,\mathrm{e}^{-cf}\,\mathrm{d}V$
$\quad-\frac{\mathrm{e}^{-cr}}{2c}\int_{\Omega_{r}}|\nabla\mathrm{R}|^{2}\,\mathrm{d}V+\frac{\mathrm{e}^{-cr}}{2}\int_{\Omega_{r}}|\nabla\mathrm{R}|^{2}\,\mathrm{d}V$
$=\int_{\Omega_{r}}\mathrm{R}\,\mathrm{Rc}(\nabla f,\nabla f)\,\mathrm{d}V_{c}.$
So the weighted integral of $\mathrm{U}_{ij}$ becomes
$\int_{\Omega_{r}}\mathrm{U}(\nabla f,\nabla f)\,\mathrm{d}V_{c}=\int_{\Omega_{r}}(\mathrm{R}^{2}-2|\mathrm{Rc}|^{2})|\nabla f|^{2}\,\mathrm{d}V_{c}.$
Now we can transform the integral using our Bochner formula from Lemma 7.2 as well as
$\int_{\Omega_{r}}$ $\mathrm{U}(\nabla f,\nabla f)\,\mathrm{d}V_{c}$
$=\frac{1}{4}\int_{\Omega_{r}}\Big(\mathrm{R}^{2}|\nabla f|^{2}-\mathrm{R}|\nabla f|^{2}+\mathrm{R}\langle\nabla\mathrm{R},\nabla f\rangle\Big)\,\mathrm{d}V_{c}$
$=\frac{1}{4}\int_{\Omega_{r}}\Big(\mathrm{R}^{2}|\nabla f|^{2}-\mathrm{R}|\nabla f|^{2}+\frac{1}{c}|\nabla\mathrm{R}|^{2}\Big)\,\mathrm{d}V_{c}-\frac{\mathrm{e}^{-cr}}{4c}\int_{\Omega_{r}}|\nabla\mathrm{R}|^{2}\,\mathrm{d}V.$
The issue is that $\mathrm{R}^{2}-2\,|\mathrm{Rc}|^{2}$ can change sign over $M$, so
$\int_{M}\big(\mathrm{R}^{2}-2\,|\mathrm{Rc}|^{2}\big)\,|\nabla f|^{2}\,\mathrm{d}V_{c}=0$
does not yield any obvious rigidity. The $\mathrm{R}^{2}|\nabla f|^{2}$ term reminds us of our work on the $\mathrm{V}$-tensor, but we must first reformulate our integral identity for $\mathrm{V}_{ij}$ for the altered-weighted volume element.
Lemma 7.4.
Let $(M^{4},f,g)$ be a complete gradient-shrinking Ricci soliton satisfying (1). Let $r>0$ be a regular value of $f$ and $\Omega_{r}=\{f\leq r\}$. Then if $\Delta\mathrm{R}=0$ and $c>0$, we have
$\int_{\Omega_{r}}\mathrm{V}(\nabla f,\nabla f)\,\mathrm{d}V_{c}=-\int_{\Omega_{r}}\Big(\frac{1}{4}\mathrm{R}^{2}|\nabla f|^{2}+\frac{3(c-1)}{2c}|\nabla\mathrm{R}|^{2}\Big)\,\mathrm{d}V_{c}+\frac{(4c-3)\mathrm{e}^{-cr}}{2c}\int_{\Omega_{r}}|\nabla\mathrm{R}|^{2}\,\mathrm{d}V$
and
$\int_{M}\mathrm{V}(\nabla f,\nabla f)\,\mathrm{d}V_{c}=-\int_{M}\Big(\frac{1}{4}\mathrm{R}^{2}|\nabla f|^{2}+\frac{3(c-1)}{2c}|\nabla\mathrm{R}|^{2}\Big)\,\mathrm{d}V_{c}.$
Proof.
We will once again integrate the terms of $\mathrm{V}(\nabla f,\nabla f)$ separately. First, we compute the following divergence:
$\nabla_{i}$ $(\nabla_{j}\mathrm{R}\nabla_{i}f\nabla_{j}f\,\mathrm{e}^{-cf})$
$=\nabla_{i}\nabla_{j}\mathrm{R}\nabla_{i}f\nabla_{j}f\mathrm{e}^{-cf}+\Delta f\langle\nabla\mathrm{R},\nabla f\rangle\mathrm{e}^{-cf}$
$\quad+\nabla^{2}f(\nabla f,\nabla\mathrm{R})\mathrm{e}^{-cf}-c|\nabla f|^{2}\langle\nabla\mathrm{R},\nabla f\rangle\mathrm{e}^{-cf}.$
Hence, the integral of $\nabla^{2}\mathrm{R}(\nabla f,\nabla f)\,\mathrm{e}^{-cf}$ is:
$\int_{\Omega_{r}}$ $\nabla^{2}\mathrm{R}(\nabla f,\nabla f)\,\mathrm{e}^{-cf}\,\mathrm{d}V$
$=\int_{\Omega_{r}}\nabla_{i}(\langle\nabla\mathrm{R},\nabla f\rangle\nabla_{i}f\,\mathrm{e}^{-cf})\,\mathrm{d}V+\int_{\Omega_{r}}(\mathrm{R}-2)\langle\nabla\mathrm{R},\nabla f\rangle\,\mathrm{e}^{-cf}\,\mathrm{d}V$
$\quad-\int_{\Omega_{r}}\Big(\frac{1}{2}\langle\nabla\mathrm{R},\nabla f\rangle-\frac{1}{2}|\nabla\mathrm{R}|^{2}\Big)\mathrm{e}^{-cf}\,\mathrm{d}V+c\int_{\Omega_{r}}(f-\mathrm{R})\langle\nabla\mathrm{R},\nabla f\rangle\mathrm{e}^{-cf}\,\mathrm{d}V$
$=\int_{\partial\Omega_{r}}|\nabla f|\langle\nabla\mathrm{R},\nabla f\rangle\mathrm{e}^{-cf}\,\mathrm{d}V$
$\quad+(c-1)\int_{\Omega_{r}}\mathrm{R}\langle\nabla\mathrm{R},\nabla f\rangle\mathrm{e}^{-cf}\,\mathrm{d}V+\frac{1}{2}\int_{\Omega_{r}}|\nabla\mathrm{R}|^{2}\mathrm{e}^{-cf}\,\mathrm{d}V$
$=-\mathrm{e}^{-cr}\Big(\frac{c-1}{c}+1\Big)\int_{\Omega_{r}}|\nabla\mathrm{R}|^{2}\,\mathrm{d}V+\Big(\frac{c-1}{c}+\frac{1}{2}\Big)\int_{\Omega_{r}}|\nabla\mathrm{R}|^{2}\mathrm{e}^{-cf}\,\mathrm{d}V.$
Therefore, we have:
$\int_{\Omega_{r}}$ $\mathrm{V}(\nabla f,\nabla f)\,\mathrm{e}^{-cf}\,\mathrm{d}V$
$=\int_{\Omega_{r}}\nabla^{2}\mathrm{R}(\nabla f,\nabla f)\,\mathrm{e}^{-cf}\,\mathrm{d}V+\frac{1}{2}\int_{\Omega_{r}}\mathrm{R}\langle\nabla\mathrm{R},\nabla f\rangle\,\mathrm{e}^{-cf}\,\mathrm{d}V$
$\quad-\frac{1}{4}\int_{\Omega_{r}}\mathrm{R}^{2}|\nabla f|^{2}\,\mathrm{e}^{-cf}\,\mathrm{d}V$
$=\mathrm{e}^{-cr}\Big(\frac{c-1}{c}+1-\frac{1}{2c}\Big)\int_{\Omega_{r}}|\nabla\mathrm{R}|^{2}\,\mathrm{d}V-\Big(\frac{c-1}{c}+\frac{1}{2}-\frac{1}{2c}\Big)\int_{\Omega_{r}}|\nabla\mathrm{R}|^{2}\,\mathrm{e}^{-cf}\,\mathrm{d}V$
$\quad-\frac{1}{4}\int_{\Omega_{r}}\mathrm{R}^{2}|\nabla f|^{2}\,\mathrm{e}^{-cf}\,\mathrm{d}V$
$=-\int_{\Omega_{r}}\Big(\frac{1}{4}\mathrm{R}^{2}|\nabla f|^{2}+\frac{3(c-1)}{2c}|\nabla\mathrm{R}|^{2}\Big)\,\mathrm{e}^{-cf}\,\mathrm{d}V+\frac{(4c-3)\mathrm{e}^{-cr}}{2c}\int_{\Omega_{r}}|\nabla\mathrm{R}|^{2}\,\mathrm{d}V.$
Now taking the limit $r\rightarrow\infty$:
$\int_{M}\mathrm{V}(\nabla f,\nabla f)\,\mathrm{e}^{-cf}\,\mathrm{d}V=-\int_{M}\Big(\frac{1}{4}\mathrm{R}^{2}|\nabla f|^{2}+\frac{3(c-1)}{2c}|\nabla\mathrm{R}|^{2}\Big)\,\mathrm{e}^{-cf}\,\mathrm{d}V.$
8 Proof of Theorem 1.1(ii)
Assume $\mathrm{U}_{ij}\equiv 0$. By Lemma 2.1, $\mathrm{tr}_{g}(\mathrm{U})=-\Delta\mathrm{R}$, hence $\Delta R\equiv 0$. Fix a regular value $r\in(0,1)$ of $f$ and write $\Omega_{r}=\{x\in M:f(x)\leq r\}$. For any $c>0$, Lemma 7.3 gives
$0=\int_{\Omega_{r}}\mathrm{U}(\nabla f,\nabla f)\,\,\mathrm{d}V_{c}=\frac{1}{4}\int_{\Omega_{r}}\!\Big((\mathrm{R}^{2}-\mathrm{R})|\nabla f|^{2}+\frac{1}{c}|\nabla\mathrm{R}|^{2}\Big)\,\mathrm{d}V_{c}-\frac{\mathrm{e}^{-cr}}{4c}\int_{\Omega_{r}}\!|\nabla\mathrm{R}|^{2}\,\,\mathrm{d}V.$
Multiplying by $4c$ and rearranging yields the basic identity
$c\int_{\Omega_{r}}(\mathrm{R}^{2}-\mathrm{R})|\nabla f|^{2}\,\mathrm{e}^{-cf}\,\,\mathrm{d}V=\int_{\Omega_{r}}|\nabla\mathrm{R}|^{2}\,\big(\mathrm{e}^{-cr}-\mathrm{e}^{-cf}\big)\,\,\mathrm{d}V.$
On the fixed bounded set $\Omega_{r}$ we have $0\leq f\leq r<1$, so all integrands are smooth and dominated by an $L^{1}$ function independent of $c$. Hence we may differentiate under the integral sign. Differentiating once gives
$\int_{\Omega_{r}}(\mathrm{R}^{2}-\mathrm{R})|\nabla f|^{2}\,\mathrm{e}^{-cf}\,\,\mathrm{d}V-c\int_{\Omega_{r}}f\,(\mathrm{R}^{2}-\mathrm{R})|\nabla f|^{2}\,\mathrm{e}^{-cf}\,\,\mathrm{d}V=\int_{\Omega_{r}}\big(f\,\mathrm{e}^{-cf}-r\,\mathrm{e}^{-cr}\big)\,|\nabla\mathrm{R}|^{2}\,\,\mathrm{d}V.$
Differentiating again yields
$\int_{\Omega_{r}}\big(cf^{2}-2f\big)(\mathrm{R}^{2}-\mathrm{R})|\nabla f|^{2}\,\mathrm{e}^{-cf}\,\,\mathrm{d}V=\int_{\Omega_{r}}\big(r^{2}\,\mathrm{e}^{-cr}-f^{2}\,\mathrm{e}^{-cf}\big)\,|\nabla\mathrm{R}|^{2}\,\,\mathrm{d}V.$
An easy induction shows that for each integer $n\geq 0$,
$\int_{\Omega_{r}}\big(nf^{\,n-1}-cf^{\,n}\big)\,(\mathrm{R}^{2}-\mathrm{R})|\nabla f|^{2}\,\mathrm{e}^{-cf}\,\,\mathrm{d}V=\int_{\Omega_{r}}\big(r^{\,n}\,\mathrm{e}^{-cr}-f^{\,n}\,\mathrm{e}^{-cf}\big)\,|\nabla\mathrm{R}|^{2}\,\,\mathrm{d}V.$ (8)
Setting $c=0$ in (8) and summing over $n\geq 0$, we use the geometric series sum formulas:
$\sum_{n=0}^{\infty}nf^{\,n-1}=\frac{1}{(1-f)^{2}},\qquad\sum_{n=0}^{\infty}f^{\,n}=\frac{1}{1-f},\qquad\sum_{n=0}^{\infty}r^{\,n}=\frac{1}{1-r}.$
This leads to the identity
$\int_{\Omega_{r}}\Bigg(\frac{(\mathrm{R}^{2}-\mathrm{R})|\nabla f|^{2}}{(1-f)^{2}}+\frac{f-r}{(1-f)(1-r)}\,|\nabla\mathrm{R}|^{2}\Bigg)\,\,\mathrm{d}V=0.$ (9)
On $\Omega_{r}$ we have $0\leq\mathrm{R}\leq f<1$, hence $\mathrm{R}^{2}-\mathrm{R}\leq 0$ and $f-r\leq 0$, while (1f)>0(1-f)>0 and (1r)>0(1-r)>0. Therefore both terms in the integrand of (9) are non-positive. Since the integral vanishes, both terms vanish identically on $\Omega_{r}$:
[…]
In particular $\nabla\mathrm{R}\equiv 0$ on $\Omega_{r}$. Together with $\Delta\mathrm{R}\equiv 0$, unique continuation implies $\mathrm{R}$ is constant on all of $M$. If $\mathrm{R}=0$ at some point then $\mathrm{R}\equiv 0$ everywhere and the soliton is isometric to the Gaussian soliton on $\mathbb{R}^{4}$. Otherwise $\mathrm{R}>0$ and the soliton is Einstein.
9 Proof of Theorem 1.2 Outside the Cone
Assume $3\beta\neq\alpha$, then $\mathrm{tr}_{g}\mathfrak{B}_{ij}=(3\beta-\alpha)\Delta\mathrm{R}=0$, hence $\Delta\mathrm{R}\equiv 0$. For any $c>0$ and a fixed regular value $r\in(0,1)$ for $f$, let $\Omega_{r}=\{f\leq r\}.$ We express the Bach-like tensor as
$\mathfrak{B}_{ij}=6\beta\,\mathrm{B}_{ij}+(\alpha-3\beta)\,\mathrm{U}_{ij}$
One can easily check that for the volume element $\mathrm{d}V_{c}$, the integration of the Bach tensor yields:
$\int_{\Omega_{r}}\mathrm{B}(\nabla f,\nabla f)\,\,\mathrm{d}V_{c}=-\frac{c}{2}\int_{\Omega_{r}}|\mathrm{D}_{ijk}|^{2}\,\mathrm{d}V_{c}=\mathrm{F}(c).$
So differentiating past the integral symbol yields
$\frac{d^{n}}{dc^{n}}\Big|_{c=0}\mathrm{F}(c)=-\frac{1}{2}\int_{\Omega_{r}}nf^{n-1}\,|\mathrm{D}_{ijk}|^{2}\,\mathrm{d}V.$
We will need to multiply both sides of the Bach-like equation by $c$ to proceed.
$\frac{d^{n}}{dc^{n}}\Big(c\,\mathrm{F}(c)\Big)=n\,\mathrm{F}^{(n-1)}(c)+c\,\mathrm{F}^{(n)}(c),$
hence
$\frac{d^{n}}{dc^{n}}\Big|_{c=0}=n\,\mathrm{F}^{(n-1)}(0)=-\frac{1}{2}\int_{\Omega_{r}}n(n-1)\,f^{n-2}\,|\mathrm{D}_{ijk}|^{2}\,\mathrm{d}V$
Summing over $n$ on the right-hand side, we obtain
$\int_{\Omega_{r}}\frac{|\mathrm{D}_{ijk}|^{2}}{(1-f)^{3}}\,\mathrm{d}V.$ (10)
Now assuming $\mathfrak{B}_{ij}\equiv 0$, and multiplying both sides by $c$:
$0=6\beta\,c\int_{\Omega_{r}}\mathrm{B}(\nabla f,\nabla f)\,\mathrm{d}V_{c}+(\alpha-3\beta)\,c\int_{\Omega_{r}}\mathrm{U}(\nabla f,\nabla f)\,\mathrm{d}V_{c},$
hence
$0=\frac{-6\beta c^{2}}{2}\int_{\Omega_{r}}|\mathrm{D}_{ijk}|^{2}\,\mathrm{d}V_{c}+(\alpha-3\beta)\,c\int_{\Omega_{r}}\mathrm{U}(\nabla f,\nabla f)\,\mathrm{d}V_{c}.$
Now differentiating with respect to $c$, evaluating each derivative at $c=0$ and summing over $n$ we finally obtain
$\boxed{-6\beta\int_{\Omega_{r}}\frac{|\mathrm{D}_{ijk}|^{2}}{(1-f)^{3}}\,\mathrm{d}V+\frac{1}{4}(\alpha-3\beta)\int_{\Omega_{r}}\Bigg(\frac{(\mathrm{R}^{2}-\mathrm{R})|\nabla f|^{2}}{(1-f)^{2}}+\frac{f-r}{(1-f)(1-r)}\,|\nabla\mathrm{R}|^{2}\Bigg)\,\,\mathrm{d}V=0.}$
Now as $0<r<1$, both terms in the second integrand are negative as in (9). If $\beta>0$ and $\alpha-3\beta>0$, or if $\beta<0$ and $\alpha-3\beta<0$, then
$|\mathrm{D}_{ijk}|^{2}\equiv 0,\quad\mathrm{R}^{2}-\mathrm{R}\equiv 0,\quad\mathrm{and}\quad|\nabla\mathrm{R}|^{2}\equiv 0\quad\mathrm{on}\,\,\Omega_{r}.$
Hence, $M$ must be Einstein or isometric to the Gaussian shrinking soliton. These ranges for $\alpha$ and $\beta$ give us the desired result for the region outside the cone $\mathcal{C}$, hence for all $(\alpha,\beta)$ aside from $(0,0)$ and the Bach-line.
10 Variational Origins of U and V
We now connect the tensors U and V with the variational structure of the Riemannian metric. In this section, we identify Bach-like tensors as the Euler–Lagrange gradients of natural quadratic curvature functionals. Throughout, we work in the standard transverse–traceless (TT) gauge, imposing for a variation $h$
$\mathrm{(TT)\,\,gauge:}\quad\mathrm{Div}_{g}(h)=0\quad\mathrm{and}\quad\mathrm{tr}_{g}(h)=0.$
Lemma 10.1.
In dimension $4$, define
$\mathcal{F}_{\alpha,\beta}(g):=-\,\alpha\int_{M}|W_{g}|^{2}\,dV_{g}\;+\;\frac{\alpha/3-\beta}{2}\int_{M}R_{g}^{2}\,dV_{g}\,.$
Then the $L^{2}$–gradient satisfies
$\mathrm{grad}\,\mathcal{F}_{\alpha,\beta}\;=\;\alpha\,\mathrm{U}\;+\;\beta\,\mathrm{V}.$
In particular, the Euler–Lagrange equation for $\mathcal{F}_{\alpha,\beta}$ is equivalent to the vanishing of the Bach-like tensor $\mathfrak{B}_{ij}=\alpha\,\mathrm{U}_{ij}+\beta\,\mathrm{V}_{ij}=0$.
Proof.
In dimension four, one has
$\delta\!\int|\mathrm{W}|^{2}\,\mathrm{d}V_{g}=-2\!\int\langle\mathrm{B},\delta g\rangle\,\mathrm{d}V_{g}\quad\mathrm{and}\quad\delta\!\int\mathrm{R}^{2}\,\mathrm{d}V_{g}=-2\!\int\langle\mathrm{V},\delta g\rangle\,\mathrm{d}V_{g}.$
Using $\mathrm{B}=\tfrac{1}{2}\,\mathrm{U}+\tfrac{1}{6}\,\mathrm{V}$,
$\delta\mathcal{F}_{\alpha,\beta}=\int\Big\langle\alpha\!\left(\tfrac{1}{2}\mathrm{U}+\tfrac{1}{6}\mathrm{V}\right)\;-\;\tfrac{\alpha/3-\beta}{2}\cdot 2\mathrm{V},\,\delta g\Big\rangle\,\mathrm{d}V_{g}=\int\langle\alpha\mathrm{U}+\beta\mathrm{V},\,\delta g\rangle\,\mathrm{d}V_{g}.$
Hence $\mathrm{grad}\,\mathcal{F}_{\alpha,\beta}=\alpha\,\mathrm{U}+\beta\,\mathrm{V}$. ∎
Corollary 10.2.
$\mathrm{V}\;=\;-\tfrac{1}{2}\,\mathrm{grad}\!\left(\int_{M}\mathrm{R}^{2}\,\,\mathrm{d}V_{g}\right),\qquad\mathrm{U}\;=\;-\,\mathrm{grad}\!\left(\int_{M}\big(|\mathrm{W}|^{2}-\tfrac{1}{6}\mathrm{R}^{2}\big)\,\,\mathrm{d}V_{g}\right).$
Proposition 10.3 (Second variation of $\int R^{2}$ in TT gauge).
Let $(M^{4},g)$ be closed. For a variation $g_{t}=g+t\,h+o(t)$ with
$\mathrm{div}h=0,\qquad\mathrm{tr}_{g}h=0\quad\text{\emph{(TT gauge)}},$
the second variation of $F(g):=\int_{M}\mathrm{R}^{2}\,\,\mathrm{d}V_{g}$ is
$\delta^{2}F[h,h]=\int_{M}\Big[-\,\mathrm{R}\,|\nabla h|^{2}\;-\;2\mathrm{R}\,\mathrm{R}_{ikjl}\,h_{ij}h_{kl}\;+\;2\mathrm{R}\,\mathrm{R}_{ik}\,h_{ij}h_{kj}\;+\;2\langle\mathrm{Rc},h\rangle^{2}\;+\;\Big(\frac{1}{2}\mathrm{R}^{2}-2\,\Delta\mathrm{R}\Big)|h|^{2}\Big]\,dV\,.\;$
Proof.
Write the first variation as $\delta F=-2\int\langle\mathrm{V},h\rangle\,\mathrm{d}V_{g}$. Then by differentiating once more, and using $\mathrm{div}h=0$, $\mathrm{tr}_{g}h=0$, the Hessian is
$\delta^{2}F[h,h]=-2\int_{M}\langle\mathrm{V}^{\prime}(h),h\rangle\,\mathrm{d}V_{g}.$
We have the following standard TT-gauge variations:
$\delta\mathrm{R}$ $=-\langle\mathrm{Rc},h\rangle,$
$\delta(\nabla_{i}\nabla_{j}\mathrm{R})$ $=\nabla_{i}\nabla_{j}(\delta\mathrm{R})-\tfrac{1}{2}(\nabla_{i}h_{j}{}^{k}+\nabla_{j}h_{i}{}^{k}-\nabla^{k}h_{ij})\,\nabla_{k}\mathrm{R},$
$\delta(\Delta\mathrm{R})$ $=\Delta(\delta\mathrm{R})-\langle h,\nabla^{2}\mathrm{R}\rangle,$
$\delta\mathrm{Rc}$ $=\tfrac{1}{2}(\nabla^{*}\nabla h)_{ij}+(\mathrm{Rm}\!*h)_{ij}-(\mathrm{Rc}\circ h+h\circ\mathrm{Rc})_{ij}.$
Using these, the Hessian can be written as
$\delta^{2}F[h,h]=\int_{M}\Big(-\mathrm{R}\,\langle\Delta_{L}h,h\rangle+2\langle\mathrm{Rc},h\rangle^{2}+(\frac{1}{2}\mathrm{R}^{2}-2\Delta\mathrm{R})|h|^{2}\Big)\,\,\mathrm{d}V_{g}.$
We expand the Lichnerowicz laplacian as $\Delta_{L}h=\nabla^{*}\nabla h+2\mathrm{Rm}\!*h-\mathrm{Rc}\circ h-h\circ\mathrm{Rc}$. Integration by parts yields the result. ∎
Remark 10.4.
If $(M^{4},g)$ is Einstein, the main contributions to the second variation come from the derivative term $\mathrm{R}\,|\nabla h|^{2}$ and the curvature terms involving $\mathrm{R}_{ikjl}\,h_{ij}\,h_{kl}$.
Proposition 10.5 (Second variation at Einstein backgrounds).
On an Einstein $4$-manifold $(M,g_{E})$ with $\mathrm{Rc}=\lambda g_{E}$ and scalar curvature $R=4\lambda$, the second variation of $\mathcal{F}_{\alpha,\beta}(g)$ in the TT-gauge is
$\delta^{2}\mathcal{F}_{\alpha,\beta}[h,h]=\int_{M}\!\Big\langle\big(\alpha\,\Delta_{L}^{2}+\frac{1}{2}\mathrm{R}\,(\beta-\tfrac{4}{3}\alpha)\,\Delta_{L}+\mathrm{R}^{2}(\tfrac{5}{36}\alpha-\tfrac{1}{4}\beta)\big)h,\;h\Big\rangle dV.$
Equivalently, for each TT eigenmode $\Delta_{L}h_{k}=\mu_{k}h_{k}$,
$\delta^{2}\mathcal{F}_{\alpha,\beta}[h,h]=\sum_{k}p_{\alpha,\beta}(\mu_{k})\,\|h_{k}\|_{L^{2}}^{2},\quad p_{\alpha,\beta}(\mu)=\alpha\mu^{2}+\frac{1}{2}\mathrm{R}\,(\beta-\tfrac{4}{3}\alpha)\mu+\mathrm{R}^{2}(\tfrac{5}{36}\alpha-\tfrac{1}{4}\beta).$
Proof.
In TT gauge at an Einstein metric, the standard linearizations give
$\mathrm{B}^{\prime}(h)=\tfrac{1}{2}\big(\Delta_{L}-\tfrac{\mathrm{R}}{3}\big)\big(\Delta_{L}-\tfrac{\mathrm{R}}{6}\big)h,\qquad\mathrm{V}^{\prime}(h)=\frac{1}{2}\mathrm{R}\,\Delta_{L}h.$
Since
$\delta^{2}\!\int_{M}|\mathrm{W}|^{2}\,\mathrm{d}V_{g}=-2\int_{M}\!\langle\mathrm{B}^{\prime}(h),h\rangle\,\mathrm{d}V_{g}\quad\mathrm{and}\quad\delta^{2}\!\int_{M}\mathrm{R}^{2}\,\mathrm{d}V_{g}=-2\int_{M}\!\langle\mathrm{V}^{\prime}(h),h\rangle\,\mathrm{d}V_{g},$
the claimed polynomial follows by expanding
$\delta^{2}\mathcal{F}_{\alpha,\beta}[h,h]=-2\alpha\!\int_{M}\langle\mathrm{B}^{\prime}(h),h\rangle\,\mathrm{d}V_{g}\;-\;(\alpha/3-\beta)\!\int_{M}\langle\mathrm{V}^{\prime}(h),h\rangle\,\mathrm{d}V_{g}.$
Corollary 10.6 (Bach line and factorization).
On the Bach line $\beta=\alpha/3$, one has
$\delta^{2}\mathcal{F}_{\alpha,\beta}[h,h]=\alpha\int_{M}\left\langle\big(\Delta_{L}-\tfrac{\mathrm{R}}{3}\big)\big(\Delta_{L}-\tfrac{\mathrm{R}}{6}\big)h,\;h\right\rangle\,\mathrm{d}V_{g},$
i.e. the Hessian factorizes into two shifted Lichnerowicz factors.
Corollary 10.7 (Local rigidity by spectral positivity).
Let $(M^{4},g_{E})$ be Einstein with TT–spectrum of $\Delta_{L}$ contained in $\mu_{0},\infty)$. If $p_{\alpha,\beta}(\mu)>0$ for every TT–eigenvalue $\mu$ (equivalently, $\inf_{\mu\geq\mu_{0}}p_{\alpha,\beta}(\mu)>0$), then $g_{E}$ is a strict local minimizer of $\mathcal{F}_{\alpha,\beta}$ modulo diffeomorphisms and scaling. In particular, there exists a neighborhood of $g_{E}$ in which the equation $\alpha U+\beta V=0$ has no nontrivial solutions other than metrics isometric to $g_{E}$.
Remark 10.8 (Flat/Gaussian background).
If $R=0=W$ (Euclidean space, or a flat compact quotient), then $B^{\prime}(h)=\tfrac{1}{2}\Delta_{L}^{2}h$ and $V^{\prime}(h)\equiv 0$ on TT. Hence
$\delta^{2}\mathcal{F}_{\alpha,\beta}[h,h]=\alpha\int_{M}\langle\Delta_{L}^{2}h,h\rangle\,dV=\alpha\int_{M}|\Delta h|^{2}\,dV\quad(\text{TT}),$
which is strictly positive for $\alpha>0$ on compactly supported (or suitably decaying/weighted) variations.

### v2

@@
 10 Variational Origins of U and V
-We now connect the tensors U and V with the variational structure of the Riemannian metric.
+The classical Bach tensor arises naturally in a number of ways.
+One approach is to start with the conformal Einstein equations and then study integrability conditions via divergence operations.
+A necessary condition to be conformally Einstein is for the Bach tensor $\mathrm{B}_{ij}$ to vanish.
+The approach taken by R.
+Bach’s [bach:1921] was to consider the first variation of the quadratic Riemannian functional
+$\mathrm{S}_{W}=\int|\mathrm{W}|^{2}\,\,\mathrm{d}V_{g}.$
+As can be seen in A. Besse [besse:1987], one has
+$\mathrm{B}\propto\mathrm{grad}\left(\int|\mathrm{W}|^{2}\,\mathrm{d}V_{g}\right),$
+so the Bach tensor is a critical point of a Riemannian functional whose energy depends on how much a metric deviates from being conformally flat.
 In this section, we identify Bach-like tensors as the Euler–Lagrange gradients of natural quadratic curvature functionals.
-Throughout, we work in the standard transverse–traceless (TT) gauge, imposing for a variation $h$
+We will work in the standard transverse–traceless (TT) gauge for the second variation, imposing for a variation $h$
 $\mathrm{(TT)\,\,gauge:}\quad\mathrm{Div}_{g}(h)=0\quad\mathrm{and}\quad\mathrm{tr}_{g}(h)=0.$
+It is worth recalling the Lichnerowicz Laplacian as it will appear in our second variation formulas. This operator $\Delta_{L}$ acts on symmetric 2-tensors via the rule
+$(\Delta_{L}\,h)_{ij}=(\nabla^{*}\nabla\,h)_{ij}+2\,\mathrm{R}_{ikjl}\,h_{kl}-\mathrm{R}_{ik}\,h_{kj}-\mathrm{R}_{jk}\,h_{ki}.$
+Here $\nabla^{*}$ denotes the divergence operator or formal adjoint of the covariant derivative $\nabla$, and $\nabla^{*}\nabla$ is the rough Laplacian.
+Note that sign conventions for the Laplacian vary, but we will take the convention that
+$\nabla^{*}\nabla\,h=-\nabla^{k}\nabla_{k}\,h.$
 Lemma 10.1.
 In dimension $4$, define
-$\mathcal{F}_{\alpha,\beta}(g):=-\,\alpha\int_{M}|W_{g}|^{2}\,dV_{g}\;+\;\frac{\alpha/3-\beta}{2}\int_{M}R_{g}^{2}\,dV_{g}\,.$
+$\mathcal{F}_{\alpha,\beta}(g):=-\,\alpha\int_{M}|\mathrm{W}_{g}|^{2}\,\mathrm{d}V_{g}\;+\;\frac{\alpha/3-\beta}{2}\int_{M}\mathrm{R}_{g}^{2}\,\mathrm{d}V_{g}\,.$
 Then the $L^{2}$–gradient satisfies
@@
 $\delta\!\int|\mathrm{W}|^{2}\,\mathrm{d}V_{g}=-2\!\int\langle\mathrm{B},\delta g\rangle\,\mathrm{d}V_{g}\quad\mathrm{and}\quad\delta\!\int\mathrm{R}^{2}\,\mathrm{d}V_{g}=-2\!\int\langle\mathrm{V},\delta g\rangle\,\mathrm{d}V_{g}.$
-Using $\mathrm{B}=\tfrac{1}{2}\,\mathrm{U}+\tfrac{1}{6}\,\mathrm{V}$,
+These standard variational identities for quadratic functionals are well-established [e.g., [besse:1987]]. Using $\mathrm{B}=\tfrac{1}{2}\,\mathrm{U}+\tfrac{1}{6}\,\mathrm{V}$,
 $\delta\mathcal{F}_{\alpha,\beta}=\int\Big\langle\alpha\!\left(\tfrac{1}{2}\mathrm{U}+\tfrac{1}{6}\mathrm{V}\right)\;-\;\tfrac{\alpha/3-\beta}{2}\cdot 2\mathrm{V},\,\delta g\Big\rangle\,\mathrm{d}V_{g}=\int\langle\alpha\mathrm{U}+\beta\mathrm{V},\,\delta g\rangle\,\mathrm{d}V_{g}.$
@@
 the second variation of $F(g):=\int_{M}\mathrm{R}^{2}\,\,\mathrm{d}V_{g}$ is
-$\delta^{2}F[h,h]=\int_{M}\Big[-\,\mathrm{R}\,|\nabla h|^{2}\;-\;2\mathrm{R}\,\mathrm{R}_{ikjl}\,h_{ij}h_{kl}\;+\;2\mathrm{R}\,\mathrm{R}_{ik}\,h_{ij}h_{kj}\;+\;2\langle\mathrm{Rc},h\rangle^{2}\;+\;\Big(\frac{1}{2}\mathrm{R}^{2}-2\,\Delta\mathrm{R}\Big)|h|^{2}\Big]\,dV\,.\;$
+$\delta^{2}F[h,h]=\int_{M}\Big[-\,\mathrm{R}\,|\nabla h|^{2}\;-\;2\mathrm{R}\,\mathrm{R}_{ikjl}\,h_{ij}h_{kl}\;+\;2\mathrm{R}\,\mathrm{R}_{ik}\,h_{ij}h_{kj}\;+\;2\langle\mathrm{Rc},h\rangle^{2}\;+\;\Big(\frac{1}{2}\mathrm{R}^{2}-2\,\Delta\mathrm{R}\Big)|h|^{2}\Big]\,\,\mathrm{d}V\,.\;$
 Proof.
@@
 On an Einstein $4$-manifold $(M,g_{E})$ with $\mathrm{Rc}=\lambda g_{E}$ and scalar curvature $R=4\lambda$, the second variation of $\mathcal{F}_{\alpha,\beta}(g)$ in the TT-gauge is
-$\delta^{2}\mathcal{F}_{\alpha,\beta}[h,h]=\int_{M}\!\Big\langle\big(\alpha\,\Delta_{L}^{2}+\frac{1}{2}\mathrm{R}\,(\beta-\tfrac{4}{3}\alpha)\,\Delta_{L}+\mathrm{R}^{2}(\tfrac{5}{36}\alpha-\tfrac{1}{4}\beta)\big)h,\;h\Big\rangle dV.$
+$\delta^{2}\mathcal{F}_{\alpha,\beta}[h,h]=\int_{M}\!\Big\langle\big(\alpha\,\Delta_{L}^{2}+\frac{1}{2}\mathrm{R}\,(\beta-\tfrac{4}{3}\alpha)\,\Delta_{L}+\mathrm{R}^{2}(\tfrac{5}{36}\alpha-\tfrac{1}{4}\beta)\big)h,\;h\Big\rangle\,\mathrm{d}V.$
 Equivalently, for each TT eigenmode $\Delta_{L}h_{k}=\mu_{k}h_{k}$,
@@
 If $R=0=W$ (Euclidean space, or a flat compact quotient), then $B^{\prime}(h)=\tfrac{1}{2}\Delta_{L}^{2}h$ and $V^{\prime}(h)\equiv 0$ on TT. Hence
-$\delta^{2}\mathcal{F}_{\alpha,\beta}[h,h]=\alpha\int_{M}\langle\Delta_{L}^{2}h,h\rangle\,dV=\alpha\int_{M}|\Delta h|^{2}\,dV\quad(\text{TT}),$
+$\delta^{2}\mathcal{F}_{\alpha,\beta}[h,h]=\alpha\int_{M}\langle\Delta_{L}^{2}h,h\rangle\,\,\mathrm{d}V=\alpha\int_{M}|\Delta h|^{2}\,\,\mathrm{d}V\quad(\text{TT}),$
 which is strictly positive for $\alpha>0$ on compactly supported (or suitably decaying/weighted) variations.

### v3

@@
-Rigidity of Gradient Shrinking Ricci Solitons with Bach-like Flatness and Related Variational Formulas
+arXiv:2511.05774v3[math.DG] 02 Mar 2026
 Abstract
-The classical Bach tensor in four dimensions can be expressed as a linear combination of two independent, symmetric, divergence-free, quadratic in curvature tensors U and V.
+The classical Bach tensor in four dimensions can be expressed as a linear combination of two independent, symmetric, divergence-free, quadratic-in-curvature tensors $\mathrm{U}$ and $\mathrm{V}$.
 Several classification results for gradient-shrinking Ricci solitons have been obtained under the assumption that the Bach tensor vanishes.
-We define a Bach-like tensor to be any other linear combination of U and V.
-We prove that, except along the Bach line, the vanishing of a Bach-like tensor forces a four-dimensional complete gradient-shrinking Ricci soliton to be either Einstein or isometric to the Gaussian soliton.
+We define a Bach-like tensor to be any other linear combination of $\mathrm{U}$ and $\mathrm{V}$.
+We prove that within a certain cone of parameters $\mathcal{C}$, the vanishing of a Bach-like tensor forces a four-dimensional complete gradient-shrinking Ricci soliton to be either Einstein or isometric to the Gaussian soliton, extending the results of Cao–Chen [4].
+The special case where $\mathrm{U}=0$ forces $f_{\min}\in\{0,1,2\}$, with rigidity holding when $f_{\min}=0,2$.
+The remaining case $f_{\min}=1$ is the central open problem, with $\mathbb{R}^{2}\times\mathbb{S}^{2}(\sqrt{2})$ as the conjectured exceptional geometry.
 Finally, we show that Bach-like tensors arise as Euler–Lagrange equations of a two-parameter family of quadratic curvature functionals and compute the corresponding first and second variation formulas.
@@
 $\mathrm{Rc}+\nabla^{2}f\;=\;\rho\,g,$ (1)
-for $\rho$ a constant.
-The soliton is said to be shrinking if $\rho>0$, and it is common to scale $f$ so that $\rho=\frac{1}{2}$ for convenience.
-Such solitons arise as self-similar solutions of Hamilton’s Ricci flow, as models for type I singularities.
-Unless otherwise stated, our work will be in dimension $n=4$.
-The classical Bach tensor was introduced by R. Bach [bach:1921] to study conformal relativity. It is defined in dimension 4 by:
+for $\rho$ a constant. The soliton is said to be shrinking if $\rho>0$, and it is common to scale $f$ so that $\rho=\frac{1}{2}$ for convenience. We normalize $f$ throughout so that
+$\mathrm{R}+|\nabla f|^{2}=f,$ (2)
+following Hamilton [6]. Such solitons arise as self-similar solutions of Hamilton’s Ricci flow, as models for type I singularities. Unless otherwise stated, our work will be in dimension $n=4$.
+The classical Bach tensor was introduced by R. Bach [2] to study conformal relativity. It is defined in dimension 4 by:
 $\mathrm{B}_{ij}=\nabla^{k}\nabla^{l}\mathrm{W}_{ijkl}+\frac{1}{2}\mathrm{R}^{kl}\mathrm{W}_{ikjl},$ (3)
-where $\mathrm{W}_{ijkl}$ denotes the Weyl curvature and $R_{ij}$ the Ricci curvature.
+where $\mathrm{W}_{ijkl}$ denotes the Weyl curvature and $\mathrm{R}_{ij}$ the Ricci curvature.
 It is well-known that $\mathrm{B}_{ij}$ is conformally invariant, symmetric, trace-free, divergence-free, and quadratic in the Riemann curvature tensor.
-As seen in [bergman:2004] there are exactly two independent symmetric 22-tensors that are divergence-free and quadratic in curvature in four dimensions.
-These are the following:
+As seen in [1] there are exactly two independent symmetric $2$-tensors that are divergence-free and quadratic in curvature in four dimensions:
 $\mathrm{U}_{ij}=2\mathrm{R}_{ipjq}\mathrm{R}^{pq}+\Delta\mathrm{R}_{ij}-\frac{1}{2}|\mathrm{Rc}|^{2}g_{ij}-\mathrm{R}\,\mathrm{R}_{ij}-\frac{1}{2}\Delta\mathrm{R}\,g_{ij}+\frac{1}{4}\mathrm{R}^{2}\,g_{ij},$ (4)
 $\mathrm{V}_{ij}=-\nabla_{i}\nabla_{j}\mathrm{R}+\Delta\mathrm{R}\,g_{ij}+\mathrm{R}\,\mathrm{R}_{ij}-\frac{1}{4}\mathrm{R}^{2}\,g_{ij}.$ (5)
-One can express the Bach tensor as a combination of $U$ and $V$. In dimension 4 we have:
+One can express the Bach tensor as a combination of $\mathrm{U}$ and $\mathrm{V}$. In dimension 4:
 $\mathrm{B}_{ij}=\frac{1}{2}\mathrm{U}_{ij}+\frac{1}{6}\mathrm{V}_{ij}.$ (6)
@@
 Cao and Q.
-Chen showed in [caochen:2013] that if the Bach tensor vanishes on a complete gradient-shrinking Ricci soliton, then it must be rigid.
-Once we move beyond the Bach tensor itself, we can show that Bach-like flatness still implies rigidity, but without the index-heavy computations typical of the Bach-flat case.
-On the one hand, our results here are somewhat stronger in that only two classifications occur; on the other, this suggests that the Bach tensor behaves more like a genuine analytic obstruction, requiring more delicate analysis that ties directly into the geometry of the conformal Cotton tensor $\mathrm{D}_{ijk}$, which will be discussed later.
-Theorem 1.1 (Vanishing Basis-Tensors).
-Let $(M^{4},g,f)$ be a complete gradient-shrinking Ricci soliton satisfying (1). Then
-  1. (i)
-    $\mathrm{V}_{ij}\equiv 0$ if and only if $M^{4}$ is either Einstein or isometric to the Gaussian shrinking soliton;
-  2. (ii)
-    $\mathrm{U}_{ij}\equiv 0$ if and only if $M^{4}$ is either Einstein or isometric to the Gaussian shrinking soliton.
+Chen showed in [4] that if the Bach tensor vanishes on a complete gradient-shrinking Ricci soliton, then it must be rigid.
+Once we move beyond the Bach tensor itself, we can show that Bach-like flatness still implies rigidity within a certain parameter cone, but without the index-heavy computations typical of the Bach-flat case.
+The principal results of this paper are as follows.
+Theorem 1.1 (Vanishing V-tensor).
+Let $(M^{4},g,f)$ be a complete gradient-shrinking Ricci soliton satisfying (1) and (2).
+Then $\mathrm{V}_{ij}\equiv 0$ if and only if $M^{4}$ is either Einstein or isometric to the Gaussian shrinking soliton.
+In particular, $f_{\min}\in\{0,2\}$ correspondingly.
+Theorem 1.2 (Vanishing U-tensor).
+Let $(M^{4},g,f)$ be a complete gradient-shrinking Ricci soliton satisfying (1) and (2).
+If $\mathrm{U}_{ij}\equiv 0$ then $f_{\min}\in\{0,1,2\}$.
+If $f_{\min}=0$ then $M^{4}$ is isometric to the Gaussian shrinking soliton, and if $f_{\min}=2$ then $M^{4}$ is Einstein.
+Conversely, every Einstein or Gaussian shrinking soliton satisfies $\mathrm{U}_{ij}\equiv 0$.
+The case $f_{\min}=1$ in Theorem 1.2 is not resolved by the methods of this paper.
+At the minimum point $p$ of $f$, the tensor $\mathrm{U}$ vanishes independently of any global assumption when $\mathrm{R}(p)=1$, so the local analysis yields no information.
+By contrast, $\mathrm{V}$ cannot vanish at $p$ when $f_{\min}=1$.
+We state the following as a conjecture.
+Conjecture 1.3.
+Let $(M^{4},g,f)$ be a complete gradient-shrinking Ricci soliton satisfying (1) and (2). If $\mathrm{U}\equiv 0$ and $f_{\min}=1$, then $M^{4}$ is isometric to $\mathbb{R}^{2}\times S^{2}(\sqrt{2})$.
+The candidate space $\mathbb{R}^{2}\times S^{2}(\sqrt{2})$ is verified to satisfy the hypotheses.
+For the shrinking soliton normalization $\rho=\tfrac{1}{2}$, the soliton equation in the $S^{2}$ directions forces radius $r=\sqrt{2}$, giving $\mathrm{R}=1$ and $f_{\min}=1$ via (2).
+Indexing by $0,1$ in the $\mathbb{R}^{2}$-directions and $a,b$ on $S^{2}(\sqrt{2})$, a direct computation gives $\mathrm{U}_{00}=\mathrm{U}_{0a}=\mathrm{U}_{ab}=0$, so $\mathrm{U}\equiv 0$.
+Moreover, $\mathrm{V}_{00}=-\tfrac{1}{4}$ and $\mathrm{V}_{ab}=\tfrac{1}{4}g_{ab}$, so $\mathrm{V}\not\equiv 0$, confirming the space is neither Einstein nor Gaussian.
+This computation was suggested to us by H.-D.
+Cao.
 Theorem 1.4 (Bach-like Rigidity).
-Let $(M^{4},g,f)$ be a complete gradient-shrinking Ricci soliton satisfying (1). Assume $(\alpha,\beta)\in\mathbb{R}^{2}\setminus\{\mathbf{0}\}$ such that $\beta\neq\alpha/3$. Then if
-$\mathfrak{B}_{ij}=\alpha\,\mathrm{U}_{ij}+\beta\,\mathrm{V}_{ij}=0,$
-$M^{4}$ is either Einstein or isometric to the Gaussian shrinking soliton.
+Let $(M^{4},g,f)$ be a complete gradient-shrinking Ricci soliton satisfying (1) and (2). Assume $(\alpha,\beta)\in\mathcal{C}$ where
+$\mathcal{C}=\{(\alpha,\beta)\,\,|\,\,\alpha\geq 0\,\,\mathrm{and}\,\,\beta>\alpha/3\}\cup\{(\alpha,\beta)\,\,|\,\,\alpha\leq 0\,\,\mathrm{and}\,\,\beta<\alpha/3\},$
+and assume $3\beta\neq\alpha$.
+If $\mathfrak{B}_{ij}=\alpha\,\mathrm{U}_{ij}+\beta\,\mathrm{V}_{ij}=0$, then $M^{4}$ is either Einstein or isometric to the Gaussian shrinking soliton.
 Conversely, every Einstein or Gaussian shrinking soliton satisfies $\mathfrak{B}_{ij}=0$.
-Hence, within the class of complete gradient-shrinking Ricci solitons, $\mathfrak{B}_{ij}=0$ is equivalent to $M^{4}$ being Einstein or isometric to the Gaussian soliton.
-We will approach the proof of Theorem 1.2 by first considering the cone
-$\mathcal{C}=\{(\alpha,\beta)\,\,|\,\,\alpha\geq 0\,\,\mathrm{and}\,\,\beta>\alpha/3\}\cup\{(\alpha,\beta)\,\,|\,\,\alpha\leq 0\,\,\mathrm{and}\,\,\beta<\alpha/3\},$
-and primarily using weighted-integral identities from the $\mathrm{V}_{ij}$ tensor.
-For the region outside the cone we will take a different approach using more general weighted-integral identities.
-This result covers the entire $(\alpha,\beta)$ plane aside from the Bach-line and the origin.
+If $3\beta=\alpha$, the rigidity results of Cao–Chen [4] apply.
+We approach the proof of Theorem 1.4 by primarily using weighted integral identities from the $\mathrm{V}_{ij}$ tensor inside the cone $\mathcal{C}$.
+This result covers the entire $(\alpha,\beta)$ plane aside from the Bach line, the $\beta=0$ axis, and the origin; a complete treatment of the remaining cases is contingent on Conjecture 1.3 and is deferred to future work.
 2 Integral Identities
-To proceed, we rely on a number of integral identities implied by $\Delta R=0$. The connection between harmonic scalar curvature and Bach-like tensors is rooted in the traces of U and V.
+To proceed, we rely on a number of integral identities implied by $\Delta\mathrm{R}=0$.
+The connection between harmonic scalar curvature and Bach-like tensors is rooted in the traces of $\mathrm{U}$ and $\mathrm{V}$.
 Lemma 2.1.
-Let $U_{ij}$ and $V_{ij}$ be given by (3) and (4) respectively. Then
-$\mathrm{tr}(\mathrm{U}_{ij})$ $=-\Delta R$
-$\mathrm{tr}(\mathrm{V}_{ij})$ $=3\Delta R.$
-So if $(\alpha,\beta)\in\mathcal{C}$, then $\mathfrak{B}_{ij}=0$ implies $\Delta R=0.$ We now derive the integral identities for both the unweighted and weighted volume elements.
+Let $\mathrm{U}_{ij}$ and $\mathrm{V}_{ij}$ be given by (4) and (5) respectively. Then
+$\mathrm{tr}(\mathrm{U}_{ij})$ $=-\Delta\mathrm{R},$
+$\mathrm{tr}(\mathrm{V}_{ij})$ $=3\Delta\mathrm{R}.$
+So if $(\alpha,\beta)\in\mathcal{C}$, then $\mathfrak{B}_{ij}=0$ implies $\Delta\mathrm{R}=0.$ We now derive the integral identities for both the unweighted and weighted volume elements.
 Lemma 2.2 (Integral Identities).
-Let $(M^{4},g,f)$ be a complete gradient-shrinking Ricci soliton satisfying (1) and assume that $\Delta\mathrm{R}=0$.
+Let $(M^{4},g,f)$ be a complete gradient-shrinking Ricci soliton satisfying (1) and (2), and assume that $\Delta\mathrm{R}=0$.
 Let $r$ be a regular value of $f$ and $\Omega_{r}=\{x\in M\,|\,f(x)\leq r\}$.
@@
 First, we have:
-$0=\int_{\Omega_{r}}\Delta R\,\mathrm{d}V=\int_{\partial\Omega_{r}}\frac{1}{|\nabla f|}\langle\nabla R,\nabla f\rangle\,\mathrm{d}S$
+$0=\int_{\Omega_{r}}\Delta\mathrm{R}\,\mathrm{d}V=\int_{\partial\Omega_{r}}\frac{1}{|\nabla f|}\langle\nabla\mathrm{R},\nabla f\rangle\,\mathrm{d}S$
 showing (1). Now we integrate by parts to get:
@@
 $=r\int_{\partial\Omega_{r}}\frac{1}{|\nabla f|}\langle\nabla\mathrm{R},\nabla f\rangle\,\mathrm{d}S-\int_{\Omega_{r}}f\Delta\mathrm{R}\,\mathrm{d}V.$
-As $\Delta R=0$ by assumption, and using (1) we see that both integrals must vanish, showing (2). To show (3) we integrate $R\Delta R$:
+As $\Delta\mathrm{R}=0$ by assumption, and using (1) we see that both integrals must vanish, showing (2). To show (3) we integrate $\mathrm{R}\Delta\mathrm{R}$:
 $0=\int_{\Omega_{r}}\mathrm{R}\Delta\mathrm{R}\,\mathrm{d}V=\int_{\partial\Omega_{r}}\frac{\mathrm{R}}{|\nabla f|}\langle\nabla\mathrm{R},\nabla f\rangle\,\mathrm{d}S-\int_{\Omega_{r}}|\nabla\mathrm{R}|^{2}\,\mathrm{d}V.$
-And by using the identity $\mathrm{R}=f-|\nabla f|^{2}$ from normalizing $f$ as in [hamilton:1988], we get (3). Now we have:
+And by using the normalization (2), we get (3). Now we have:
 $0$ $=\int_{\Omega_{r}}\Delta\mathrm{R}\,\mathrm{e}^{-f}\,\mathrm{d}V$
@@
 $=\int_{\Omega_{r}}\langle\nabla\mathrm{R},\nabla f\rangle\,\mathrm{e}^{-f}\,\mathrm{d}V$
-showing (4). To show (5) we integrate $R\Delta R$ with the weighted volume $\mathrm{e}^{-f}\,\mathrm{d}V$:
+showing (4). To show (5) we integrate $\mathrm{R}\Delta\mathrm{R}$ with the weighted volume $\mathrm{e}^{-f}\,\mathrm{d}V$:
 $0$ $=\int_{\Omega_{r}}\mathrm{R}\Delta\mathrm{R}\,\mathrm{e}^{-f}\,\mathrm{d}V$
 $=\int_{\partial\Omega_{r}}\frac{\mathrm{R}}{|\nabla f|}\langle\nabla\mathrm{R},\nabla f\rangle\,\mathrm{e}^{-f}\,\mathrm{d}S-\int_{\Omega_{r}}\langle\nabla\big(\mathrm{R}\,\mathrm{e}^{-f}\big),\nabla\mathrm{R}\rangle\,\mathrm{d}V$
-$=\int_{\partial\Omega_{r}}\frac{\mathrm{R}}{|\nabla f|}\langle\nabla\mathrm{R},\nabla f\rangle\,\mathrm{e}^{-f}\,\mathrm{d}S-\int_{\Omega_{r}}\Big(|\nabla\mathrm{R}|^{2}-R\langle\nabla\mathrm{R},\nabla f\rangle\Big)\,\mathrm{e}^{-f}\,\mathrm{d}V$
+$=\int_{\partial\Omega_{r}}\frac{\mathrm{R}}{|\nabla f|}\langle\nabla\mathrm{R},\nabla f\rangle\,\mathrm{e}^{-f}\,\mathrm{d}S-\int_{\Omega_{r}}\Big(|\nabla\mathrm{R}|^{2}-\mathrm{R}\langle\nabla\mathrm{R},\nabla f\rangle\Big)\,\mathrm{e}^{-f}\,\mathrm{d}V$
 $=-\int_{\partial\Omega_{r}}|\nabla f|\langle\nabla\mathrm{R},\nabla f\rangle\,\mathrm{e}^{-f}\,\mathrm{d}S-\int_{\Omega_{r}}\Big(|\nabla\mathrm{R}|^{2}-2\mathrm{R}\,\mathrm{Rc}(\nabla f,\nabla f)\Big)\,\mathrm{e}^{-f}\,\mathrm{d}V.$
@@
 We will obtain integral identities by assuming $\mathrm{V}_{ij}=0$, and use these to prove Theorem 1.1.
-We first consider the Hessian term:
+We first consider the Hessian term.
 Lemma 3.1.
-Let $(M^{4},g,f)$ be a complete gradient-shrinking Ricci soliton satisfying (1) and assume that $\Delta R=0$. Then if $r$ is a regular value of $f$ and $\Omega_{r}=\{x\in M\,|\,f(x)\leq r\}$, we have:
+Let $(M^{4},g,f)$ be a complete gradient-shrinking Ricci soliton satisfying (1) and (2), and assume that $\Delta\mathrm{R}=0$.
+Then if $r$ is a regular value of $f$ and $\Omega_{r}=\{x\in M\,|\,f(x)\leq r\}$, we have:
 $\int_{\Omega_{r}}\nabla^{2}\mathrm{R}(\nabla f,\nabla f)\,\mathrm{e}^{-f}\,\mathrm{d}V=-\mathrm{e}^{-r}\int_{\Omega_{r}}|\nabla\mathrm{R}|^{2}\,\mathrm{d}V+\frac{1}{2}\int_{\Omega_{r}}|\nabla\mathrm{R}|^{2}\,\mathrm{e}^{-f}\,\mathrm{d}V.$
@@
 By the Cauchy-Schwarz inequality we have:
-$\frac{1}{4}|\nabla\mathrm{R}|^{2}=|\mathrm{Rc}(\nabla f)|^{2}\leq|\mathrm{Rc}|^{2}|\nabla f|^{2}$
-By H.-D. Cao and D. Zhou’s growth estimates for the potential function [caozhou:2010], we have:
-$|\nabla f|^{2}(x)\leq\frac{1}{4}(d(x)+c_{2})^{2}$
+$\frac{1}{4}|\nabla\mathrm{R}|^{2}=|\mathrm{Rc}(\nabla f)|^{2}\leq|\mathrm{Rc}|^{2}|\nabla f|^{2}.$
+By H.-D. Cao and D. Zhou’s growth estimates for the potential function [5], we have:
+$|\nabla f|^{2}(x)\leq\frac{1}{4}(d(x)+c_{2})^{2},$
 […]
 $\int_{M}|\mathrm{Rc}|^{2}|\nabla f|^{2}\mathrm{e}^{-\alpha f}\,\mathrm{d}V=\int_{K}|\mathrm{Rc}|^{2}|\nabla f|^{2}\mathrm{e}^{-\alpha f}\,\mathrm{d}V+\int_{M\backslash K}|\mathrm{Rc}|^{2}|\nabla f|^{2}\mathrm{e}^{-\alpha f}\,\mathrm{d}V.$
-Clearly the first integral over the compact set $K$ is finite. Now we have for the second integral:
+The first integral over the compact set $K$ is finite. We show that the second integral is also finite:
 $\begin{split}\int_{M\backslash K}|\mathrm{Rc}|^{2}|\nabla f|^{2}\mathrm{e}^{-\alpha f}\,\mathrm{d}V&=\int_{M\backslash K}(|\mathrm{Rc}|^{2}\mathrm{e}^{-\alpha f/2})|\nabla f|^{2}\mathrm{e}^{-\alpha f/2}\,\mathrm{d}V\\ &\leq\int_{M\backslash K}|\mathrm{Rc}|^{2}\mathrm{e}^{-\alpha f/2}\,\mathrm{d}V.\end{split}$
-By O. Munteanu and N. Sesum’s results in [munteanusesum:2013], we have:
+By O. Munteanu and N. Sesum’s results in [7], we have:
 $\int_{M}|\mathrm{Rc}|^{2}\mathrm{e}^{-\lambda f}\,\mathrm{d}V<\infty,$
@@
 We first apply Lemma 3.2:
-$\int_{M}|\nabla\mathrm{R}|^{2}\mathrm{e}^{-\alpha f/2}\,\mathrm{d}V<\infty$
+$\int_{M}|\nabla\mathrm{R}|^{2}\mathrm{e}^{-\alpha f/2}\,\mathrm{d}V<\infty.$
 Since $f\leq r$ on $\Omega_{r}$, we also have $\mathrm{e}^{-\alpha r/2}\leq\mathrm{e}^{-\alpha f/2}$:
@@
 $\lim_{r\rightarrow\infty}\Big(\frac{\mathrm{e}^{-\alpha r}}{2}\int_{\Omega_{r}}|\nabla\mathrm{R}|^{2}\,\mathrm{d}V\Big)\leq\lim_{r\rightarrow\infty}\Big(\frac{\mathrm{e}^{-\alpha r/2}}{2}\int_{\Omega_{r}}|\nabla\mathrm{R}|^{2}\mathrm{e}^{-\alpha f/2}\,\mathrm{d}V\Big)=0.$
-4 Proof of Theorem 1.1(i)
+4 Proof of Theorem 1.1
 Proof.
@@
 $=-\frac{1}{2}\int_{\Omega_{r}}|\nabla\mathrm{R}|^{2}\,\mathrm{e}^{-r}\,\mathrm{d}V+\frac{1}{2}\int_{\Omega_{r}}|\nabla\mathrm{R}|^{2}\,\mathrm{e}^{-f}\,\mathrm{d}V.$
-And using Lemma 3.1, we compute the integral of $V(\nabla f,\nabla f)$ with respect to the weighted volume element $\mathrm{e}^{-f}\,\mathrm{d}V$, and we note that the second term involving $\Delta R$ again vanishes:
+And using Lemma 3.1, we compute the integral of $\mathrm{V}(\nabla f,\nabla f)$ with respect to the weighted volume element $\mathrm{e}^{-f}\,\mathrm{d}V$, noting that the term involving $\Delta\mathrm{R}$ vanishes:
 $\int_{\Omega_{r}}\mathrm{V}($ $\nabla f,\nabla f)\,\mathrm{e}^{-f}\,\mathrm{d}V$
@@
 $\int_{M}\mathrm{V}(\nabla f,\nabla f)\,\mathrm{e}^{-f}\,\mathrm{d}V=-\int_{M}\frac{1}{4}\mathrm{R}^{2}|\nabla f|^{2}\,\mathrm{e}^{-f}\,\mathrm{d}V.$ (8)
-Assuming $\mathrm{V}_{ij}=0$, for all $p\in M$ we must have either $\mathrm{R}=0$ or $|\nabla f|=0$.
+Assuming $\mathrm{V}_{ij}=0$, we must have either $\mathrm{R}=0$ or $|\nabla f|=0$ pointwise.
 By the work of S.
@@
 Rimoldi, and A.
-Setti [pigolarimoldi:2011], either $R>0$ everywhere, hence $\nabla f=0$ and the manifold is Einstein, or if $R=0$ anywhere, then the soliton must be isometric to the Gaussian soliton on $\mathbb{R}^{4}$.
+Setti [8], either $\mathrm{R}>0$ everywhere, hence $\nabla f=0$ and the manifold is Einstein, or if $\mathrm{R}=0$ anywhere, then the soliton must be isometric to the Gaussian soliton on $\mathbb{R}^{4}$.
+The values $f_{\min}\in\{0,2\}$ follow from (2): if Einstein then $\nabla f\equiv 0$ and $\mathrm{Rc}=\tfrac{1}{2}g$, giving $\mathrm{R}=2$ and $f_{\min}=2$; if Gaussian then $\mathrm{R}=0$ and $f_{\min}=0$.
 For the converse, suppose $(M^{4},g,f)$ is Einstein. Then as the scalar curvature is constant, we have:
-$\mathrm{V}_{ij}$ $=\frac{1}{4}\mathrm{R}^{2}g_{ij}-\mathrm{R}\,\mathrm{R}_{ij}$
-$=-\mathrm{R}(\mathrm{R}_{ij}-\frac{1}{4}\mathrm{R}\,g_{ij})$
-$=0.$
+$\mathrm{V}_{ij}$ $=\mathrm{R}\,\mathrm{R}_{ij}-\frac{1}{4}\mathrm{R}^{2}g_{ij}=-\mathrm{R}\Big(\mathrm{R}_{ij}-\frac{1}{4}\mathrm{R}\,g_{ij}\Big)=0.$
 If the manifold is Gaussian, then $\mathrm{R}=0$ and again $\mathrm{V}=0$. ∎
 5 Weighted Bach Integral
-In [caochen:2013] H.-D.
+In [4] H.-D.
 Cao and Q.
@@
 $\mathrm{C}_{ijk}=-\frac{n-2}{n-3}\nabla_{l}\mathrm{W}_{ijkl},$
-where $\mathrm{W}_{ijkl}$ is the Weyl curvature. The Cotton tensor is skew-symmetric in the first two indices and trace-free in any two indices. That is,
+where $\mathrm{W}_{ijkl}$ is the Weyl curvature. The Cotton tensor is skew-symmetric in the first two indices and trace-free in any two indices:
 $\mathrm{C}_{ijk}=-\mathrm{C}_{jik}\quad\mathrm{and}\quad g^{ij}\,\mathrm{C}_{ijk}=g^{ik}\,\mathrm{C}_{ijk}=0.$
@@
 On a complete gradient-shrinking Ricci soliton, $\mathrm{D}_{ijk}$ can be expressed as:
-$D_{ijk}$ $=\frac{1}{n-2}\big(\mathrm{R}_{jk}\nabla_{i}f-\mathrm{R}_{ik}\nabla_{j}f\big)+\frac{1}{2(n-1)(n-2)}\big(g_{jk}\,\nabla_{i}\mathrm{R}-g_{ik}\,\nabla_{j}\mathrm{R}\big)$
+$\mathrm{D}_{ijk}$ $=\frac{1}{n-2}\big(\mathrm{R}_{jk}\nabla_{i}f-\mathrm{R}_{ik}\nabla_{j}f\big)+\frac{1}{2(n-1)(n-2)}\big(g_{jk}\,\nabla_{i}\mathrm{R}-g_{ik}\,\nabla_{j}\mathrm{R}\big)$
 $\quad-\frac{1}{(n-1)(n-2)}\mathrm{R}\,\big(g_{jk}\,\nabla_{i}f-g_{ik}\,\nabla_{j}f\big).$
-H.-D. Cao and Q. Chen [caochen:2013] showed how the Bach tensor can be expressed in terms of $\mathrm{C}_{ijk}$ and $\mathrm{D}_{ijk}$:
+H.-D. Cao and Q. Chen [4] showed how the Bach tensor can be expressed in terms of $\mathrm{C}_{ijk}$ and $\mathrm{D}_{ijk}$:
 $\mathrm{B}_{ij}=-\frac{1}{n-2}\Big(\nabla_{k}\mathrm{D}_{ikj}+\frac{n-3}{n-2}\mathrm{C}_{jli}\nabla_{l}f\Big).$
-We can now link together the integral of the Bach tensor contracted with the gradient of the potential function to the norm of $\mathrm{D}_{ijk}$.
 Lemma 5.1.
@@
 Proof.
-Let $r$ be a regular value of $f$ and $\Omega_{r}=\Omega_{r}=\{f\leq r\}$. We use the antisymmetric properties of the Cotton and $\mathrm{D}_{ijk}$ tensor:
+Let $r$ be a regular value of $f$ and $\Omega_{r}=\{f\leq r\}$. We use the antisymmetric properties of the Cotton and $\mathrm{D}_{ijk}$ tensor:
 $(n-2)\int_{\Omega_{r}}$ $\mathrm{B}_{ij}\nabla_{i}f\nabla_{j}f\,\mathrm{e}^{-f}\,\mathrm{d}V$
@@
 6 Proof of Theorem 1.4
-First observe that for any $(\alpha,\beta)\in\mathcal{C}$, and a Bach-like tensor $\mathfrak{B}_{ij}=\alpha\,\mathrm{U}_{ij}+\beta\,\mathrm{V}_{ij}$ we can write:
+First observe that for any $(\alpha,\beta)\in\mathcal{C}$, we can write:
 $\mathfrak{B}_{ij}=2\alpha\,\mathrm{B}_{ij}+\Big(\beta-\frac{\alpha}{3}\Big)\,\mathrm{V}_{ij}.$
@@
 $\int_{M}\mathfrak{B}(\nabla f,\nabla f)\,\mathrm{e}^{-f}\,\mathrm{d}V$ $=-\alpha\int_{M}|\mathrm{D}_{ijk}|^{2}\,\mathrm{e}^{-f}\,\mathrm{d}V-\frac{1}{4}\Big(\beta-\frac{\alpha}{3}\Big)\int_{M}\mathrm{R}^{2}\,|\nabla f|^{2}\,\mathrm{e}^{-f}\,\,\mathrm{d}V=0.$
-Since $(\alpha,\beta)\in\mathcal{C}$, we have two cases.
-If $\alpha>0$ and $\beta>\alpha/3$, then we must have $|\mathrm{D}_{ijk}|^{2}(p)=0$ and $\mathrm{R}^{2}(p)|\nabla f(p)|^{2}=0$ for all $p\in M$.
-Then we can conclude that $M$ is either Einstein or isometric to the Gaussian soliton as in the proof of Theorem 1.1.
-The argument for when $\alpha<0$ and $\beta<\alpha/3$ is analogous.
+Since $(\alpha,\beta)\in\mathcal{C}$, if $\alpha>0$ and $\beta>\alpha/3$ then both integrals are nonnegative and must each vanish, so $|\mathrm{D}_{ijk}|^{2}\equiv 0$ and $\mathrm{R}^{2}|\nabla f|^{2}\equiv 0$.
+We conclude that $M$ is either Einstein or isometric to the Gaussian soliton as in the proof of Theorem 1.1.
+The argument for $\alpha<0$ and $\beta<\alpha/3$ is analogous.
 7 The Vanishing U-tensor
-In order to extend the $(\alpha,\beta)$-cone in Theorem 1.2, we need to consider the implications of $\mathrm{U}_{ij}=0$.
-We attempt the same style of argument as we did with $\mathrm{V}_{ij}=0$.
-Define a weighted volume element of the form
+The analysis we used for the vanishing $V$-tensor doesn’t completely carry over to $\mathrm{U}=0$, but will allow us to prove Theorem 1.2. Define a weighted volume element of the form
 $\mathrm{d}V_{c}(\mathbf{x})=\mathrm{e}^{-cf(\textbf{x})}\,\mathrm{d}V\quad\mathrm{where}\quad c>0.$
 Lemma 7.1.
-Let $(M^{4},f,g)$ be a complete gradient-shrinking Ricci soliton satisfying (1), normalized so that $\mathrm{R}+|\nabla f|^{2}=f$.
+Let $(M^{4},f,g)$ be a complete gradient-shrinking Ricci soliton satisfying (1) and (2).
 Let $r$ be a regular value of $f$ and define the sub-level set $\Omega_{r}=\{x\in M\,|\,f(x)\leq r\}$.
@@
   3. (3.)
-    $\int_{\Omega_{r}}f\,\langle\nabla R,\nabla f\rangle\,\mathrm{d}V_{c}=0$.
+    $\int_{\Omega_{r}}f\,\langle\nabla\mathrm{R},\nabla f\rangle\,\mathrm{d}V_{c}=0$.
 Proof.
@@
 $=\int_{\partial\Omega_{r}}\frac{1}{|\nabla f|}\langle\nabla\mathrm{R},\nabla f\rangle\mathrm{e}^{-cf}\,\mathrm{d}S+c\int_{\Omega_{r}}\langle\nabla\mathrm{R},\nabla f\rangle\mathrm{e}^{-cf}\,\,\mathrm{d}V$
-$=c\int_{\Omega_{r}}\langle\nabla\mathrm{R},\nabla f\rangle\mathrm{e}^{-f}\,\,\mathrm{d}V.$
-Next, we have:
+$=c\int_{\Omega_{r}}\langle\nabla\mathrm{R},\nabla f\rangle\mathrm{e}^{-cf}\,\,\mathrm{d}V,$
+showing (1). Next, we have:
 $0$ $=\int_{\Omega_{r}}\mathrm{R}\Delta\mathrm{R}\,\mathrm{e}^{-cf}\,\mathrm{d}V$
 $=\int_{\partial\Omega_{r}}\frac{\mathrm{R}}{|\nabla f|}\langle\nabla\mathrm{R},\nabla f\rangle\,\mathrm{e}^{-cf}\,\mathrm{d}S-\int_{\Omega_{r}}\Big(|\nabla\mathrm{R}|^{2}-c\,\mathrm{R}\,\langle\nabla\mathrm{R},\nabla f\rangle\Big)\mathrm{e}^{-cf}\,\mathrm{d}V.$
-Hence,
+Hence, using Lemma 2.2(3),
 $\int_{\Omega_{r}}\mathrm{R}\,\langle\nabla\mathrm{R},\nabla f\rangle\,\mathrm{e}^{-cf}\,\mathrm{d}V=\frac{1}{c}\int_{\Omega_{r}}|\nabla\mathrm{R}|^{2}\mathrm{e}^{-cf}\,\mathrm{d}V-\frac{\mathrm{e}^{-c\,r}}{c}\int_{\Omega_{r}}|\nabla\mathrm{R}|^{2}\,\mathrm{d}V,$
-where we used 2.2(3). And the last critical identity is:
+showing (2). And the last identity is:
 $\int_{\Omega_{r}}f$ $\langle\nabla\mathrm{R},\nabla f\rangle\mathrm{e}^{-cf}\,\mathrm{d}V$
@@
 $=\frac{1}{c}\int_{\Omega_{r}}\Big(\langle\nabla\mathrm{R},\nabla f\rangle+f\Delta\mathrm{R}\Big)\mathrm{e}^{-cf}\,\mathrm{d}V$
-$=0.$
-We first state a quick consequence of Bochner’s formula, and then we will proceed with the main integral identity.
+$=0,$
+showing (3). ∎
+We first state a quick consequence of Bochner’s formula.
 Lemma 7.2.
-Let $(M^{4},f,g)$ be a complete gradient-shrinking Ricci soliton satisfying (1), normalized so that $\mathrm{R}+|\nabla f|^{2}=f$. Then if $\Delta\mathrm{R}=0$:
-$2|\mathrm{Rc}|^{2}=\mathrm{R}+\langle\nabla\mathrm{R},\nabla f\rangle$
+Let $(M^{4},f,g)$ be a complete gradient-shrinking Ricci soliton satisfying (1) and (2). If $\Delta\mathrm{R}=0$, then:
+$2|\mathrm{Rc}|^{2}=\mathrm{R}+\langle\nabla\mathrm{R},\nabla f\rangle.$
 Proof.
-By Bochner’s formula, we have:
+By Bochner’s formula:
 $\frac{1}{2}\Delta|\nabla f|^{2}$ $=\langle\nabla(\Delta f),\nabla f\rangle+|\mbox{Hess}(f)|^{2}+\mathrm{Rc}(\nabla f,\nabla f)$
 $\frac{1}{2}\Delta(f-\mathrm{R})$ $=\langle\nabla(2-\mathrm{R}),\nabla f\rangle+|\mbox{Hess}(f)|^{2}+\frac{1}{2}\langle\nabla\mathrm{R},\nabla f\rangle$
-$\frac{1}{2}(2-\mathrm{R})$ $=-\frac{1}{2}\langle\nabla\mathrm{R},\nabla f\rangle+|\frac{1}{2}g-\mathrm{Rc}|^{2}$
+$\frac{1}{2}(2-\mathrm{R})$ $=-\frac{1}{2}\langle\nabla\mathrm{R},\nabla f\rangle+\left|\frac{1}{2}g-\mathrm{Rc}\right|^{2}$
 $1-\frac{\mathrm{R}}{2}$ $=-\frac{1}{2}\langle\nabla\mathrm{R},\nabla f\rangle+1-\mathrm{R}+|\mathrm{Rc}|^{2}$
@@
 Lemma 7.3.
-Let $(M^{4},g,f)$ be a complete gradient-shrinking Ricci soliton satisfying (1). Assume that $\Delta\mathrm{R}\equiv 0$, $r>0$ is a regular value of $f$, and $\Omega_{r}=\{f\leq r\}$. Then
+Let $(M^{4},g,f)$ be a complete gradient-shrinking Ricci soliton satisfying (1) and (2). Assume that $\Delta\mathrm{R}\equiv 0$, $r>0$ is a regular value of $f$, and $\Omega_{r}=\{f\leq r\}$. Then
 $\int_{\Omega_{r}}\mathrm{U}(\nabla f,\nabla f)\,\mathrm{d}V_{c}=\frac{1}{4}\int_{\Omega_{r}}\big(\mathrm{R}^{2}-2\,|\mathrm{Rc}|^{2}\big)\,|\nabla f|^{2}\,\mathrm{d}V_{c}.$
 Moreover, this can be rewritten as
-$\int_{\Omega_{r}}\mathrm{U}(\nabla f,\nabla f)\,\mathrm{d}V_{c}=\frac{1}{4}\int_{\Omega_{r}}\big(\mathrm{R}^{2}\,|\nabla f|^{2}-\mathrm{R}\,|\nabla f|^{2}+\frac{1}{c}|\nabla\mathrm{R}|^{2}\big)\,\mathrm{d}V_{c}-\frac{\mathrm{e}^{-cr}}{4c}\int_{\Omega_{r}}|\nabla\mathrm{R}|^{2}\,\mathrm{d}V.$
+$\int_{\Omega_{r}}\mathrm{U}(\nabla f,\nabla f)\,\mathrm{d}V_{c}=\frac{1}{4}\int_{\Omega_{r}}\big(\mathrm{R}^{2}\,|\nabla f|^{2}-\mathrm{R}\,|\nabla f|^{2}+\tfrac{1}{c}|\nabla\mathrm{R}|^{2}\big)\,\mathrm{d}V_{c}-\frac{\mathrm{e}^{-cr}}{4c}\int_{\Omega_{r}}|\nabla\mathrm{R}|^{2}\,\mathrm{d}V.$
 Proof.
 One can easily check that on a soliton we have
-$U_{ij}=\mathrm{R}_{ij}+\nabla_{k}\mathrm{R}_{ij}\nabla_{k}f-\frac{1}{2}|\mathrm{Rc}|^{2}g_{ij}-\mathrm{R}\,\mathrm{R}_{ij}-\frac{1}{2}\Delta\mathrm{R}g_{ij}+\frac{\mathrm{R}^{2}}{4}g_{ij},$
+$\mathrm{U}_{ij}=\mathrm{R}_{ij}+\nabla_{k}\mathrm{R}_{ij}\nabla_{k}f-\frac{1}{2}|\mathrm{Rc}|^{2}g_{ij}-\mathrm{R}\,\mathrm{R}_{ij}-\frac{1}{2}\Delta\mathrm{R}\,g_{ij}+\frac{\mathrm{R}^{2}}{4}g_{ij},$
 as well as the soliton identities
   1. (a)
-    $\mathrm{\mathrm{Rc}}(\nabla f)=\frac{1}{2}\nabla\mathrm{R}\implies\mathrm{\mathrm{Rc}}(\nabla f,\nabla f)=\frac{1}{2}\langle\nabla\mathrm{R},\nabla f\rangle$,
+    $\mathrm{Rc}(\nabla f)=\frac{1}{2}\nabla\mathrm{R}\implies\mathrm{Rc}(\nabla f,\nabla f)=\frac{1}{2}\langle\nabla\mathrm{R},\nabla f\rangle$,
   2. (b)
@@
     $\nabla^{2}f(\nabla f)=\frac{1}{2}\nabla f-\frac{1}{2}\nabla\mathrm{R}$.
-The second term in our expression for $U$ is simplified by looking at the following divergence:
+The second term in our expression for $\mathrm{U}$ is simplified by looking at the following divergence:
 $\nabla_{k}(\mathrm{R}_{ij}$ $\nabla_{i}f\nabla_{j}f\nabla_{k}f\,\mathrm{e}^{-cf})$
@@
 $\int_{\Omega_{r}}$ $\nabla_{k}\mathrm{R}_{ij}\nabla_{i}f\nabla_{j}f\nabla_{k}f\,\mathrm{e}^{-cf}\,\mathrm{d}V$
-$=\int_{\Omega_{r}}\mathrm{div}(\frac{1}{2}\langle\nabla\mathrm{R},\nabla f\rangle\nabla_{k}f\,\mathrm{e}^{-cf})\,\mathrm{d}V+\frac{1}{2}\int_{\Omega_{r}}\mathrm{R}\langle\nabla\mathrm{R},\nabla f\rangle\,\mathrm{e}^{-cf}\,\mathrm{d}V$
+$=\int_{\Omega_{r}}\mathrm{div}\left(\frac{1}{2}\langle\nabla\mathrm{R},\nabla f\rangle\nabla_{k}f\,\mathrm{e}^{-cf}\right)\,\mathrm{d}V+\frac{1}{2}\int_{\Omega_{r}}\mathrm{R}\langle\nabla\mathrm{R},\nabla f\rangle\,\mathrm{e}^{-cf}\,\mathrm{d}V$
 $\quad-\frac{1}{2}\int_{\Omega_{r}}|\nabla\mathrm{R}|^{2}\,\mathrm{e}^{-cf}\,\mathrm{d}V+\frac{\mathrm{e}^{-cr}}{2}\int_{\Omega_{r}}|\nabla\mathrm{R}|^{2}\,\mathrm{d}V$
@@
 So the weighted integral of $\mathrm{U}_{ij}$ becomes
-$\int_{\Omega_{r}}\mathrm{U}(\nabla f,\nabla f)\,\mathrm{d}V_{c}=\int_{\Omega_{r}}(\mathrm{R}^{2}-2|\mathrm{Rc}|^{2})|\nabla f|^{2}\,\mathrm{d}V_{c}.$
-Now we can transform the integral using our Bochner formula from Lemma 7.2 as well as
+$\int_{\Omega_{r}}\mathrm{U}(\nabla f,\nabla f)\,\mathrm{d}V_{c}=\frac{1}{4}\int_{\Omega_{r}}(\mathrm{R}^{2}-2|\mathrm{Rc}|^{2})|\nabla f|^{2}\,\mathrm{d}V_{c}.$
+Now transforming using the Bochner formula from Lemma 7.2:
 $\int_{\Omega_{r}}$ $\mathrm{U}(\nabla f,\nabla f)\,\mathrm{d}V_{c}$
@@
 $\int_{M}\big(\mathrm{R}^{2}-2\,|\mathrm{Rc}|^{2}\big)\,|\nabla f|^{2}\,\mathrm{d}V_{c}=0$
-does not yield any obvious rigidity.
-The $\mathrm{R}^{2}|\nabla f|^{2}$ term reminds us of our work on the $\mathrm{V}$-tensor, but we must first reformulate our integral identity for $\mathrm{V}_{ij}$ for the altered-weighted volume element.
+does not yield obvious rigidity.
+For comparison, we state the analogous identity for $\mathrm{V}(\nabla f,\nabla f)$ under $\mathrm{d}V_{c}$.
+The contrast with Lemma 7.3 is instructive: for $c\geq 1$ the integrand for $\mathrm{V}$ acquires a nonnegative term, while the integrand for $\mathrm{U}$ remains sign-indefinite for any $c>0$.
+This further highlights the subtleties in the $\mathrm{U}=0$ case.
 Lemma 7.4.
-Let $(M^{4},f,g)$ be a complete gradient-shrinking Ricci soliton satisfying (1). Let $r>0$ be a regular value of $f$ and $\Omega_{r}=\{f\leq r\}$. Then if $\Delta\mathrm{R}=0$ and $c>0$, we have
-$\int_{\Omega_{r}}\mathrm{V}(\nabla f,\nabla f)\,\mathrm{d}V_{c}=-\int_{\Omega_{r}}\Big(\frac{1}{4}\mathrm{R}^{2}|\nabla f|^{2}+\frac{3(c-1)}{2c}|\nabla\mathrm{R}|^{2}\Big)\,\mathrm{d}V_{c}+\frac{(4c-3)\mathrm{e}^{-cr}}{2c}\int_{\Omega_{r}}|\nabla\mathrm{R}|^{2}\,\mathrm{d}V$
-and
+Let $(M^{4},f,g)$ be a complete gradient-shrinking Ricci soliton satisfying (1) and (2). Let $r>0$ be a regular value of $f$ and $\Omega_{r}=\{f\leq r\}$. If $\Delta\mathrm{R}=0$ and $c>0$, then
+$\int_{\Omega_{r}}\mathrm{V}(\nabla f,\nabla f)\,\mathrm{d}V_{c}=-\int_{\Omega_{r}}\Big(\frac{1}{4}\mathrm{R}^{2}|\nabla f|^{2}+\frac{3(c-1)}{2c}|\nabla\mathrm{R}|^{2}\Big)\,\mathrm{d}V_{c}+\frac{(4c-3)\mathrm{e}^{-cr}}{2c}\int_{\Omega_{r}}|\nabla\mathrm{R}|^{2}\,\mathrm{d}V,$
+and in particular
 $\int_{M}\mathrm{V}(\nabla f,\nabla f)\,\mathrm{d}V_{c}=-\int_{M}\Big(\frac{1}{4}\mathrm{R}^{2}|\nabla f|^{2}+\frac{3(c-1)}{2c}|\nabla\mathrm{R}|^{2}\Big)\,\mathrm{d}V_{c}.$
 Proof.
-We will once again integrate the terms of $\mathrm{V}(\nabla f,\nabla f)$ separately. First, we compute the following divergence:
+We compute the following divergence:
 $\nabla_{i}$ $(\nabla_{j}\mathrm{R}\nabla_{i}f\nabla_{j}f\,\mathrm{e}^{-cf})$
@@
 $\quad+\nabla^{2}f(\nabla f,\nabla\mathrm{R})\mathrm{e}^{-cf}-c|\nabla f|^{2}\langle\nabla\mathrm{R},\nabla f\rangle\mathrm{e}^{-cf}.$
-Hence, the integral of $\nabla^{2}\mathrm{R}(\nabla f,\nabla f)\,\mathrm{e}^{-cf}$ is:
+Hence the integral of $\nabla^{2}\mathrm{R}(\nabla f,\nabla f)\,\mathrm{e}^{-cf}$ is:
 $\int_{\Omega_{r}}$ $\nabla^{2}\mathrm{R}(\nabla f,\nabla f)\,\mathrm{e}^{-cf}\,\mathrm{d}V$
@@
 $\quad-\int_{\Omega_{r}}\Big(\frac{1}{2}\langle\nabla\mathrm{R},\nabla f\rangle-\frac{1}{2}|\nabla\mathrm{R}|^{2}\Big)\mathrm{e}^{-cf}\,\mathrm{d}V+c\int_{\Omega_{r}}(f-\mathrm{R})\langle\nabla\mathrm{R},\nabla f\rangle\mathrm{e}^{-cf}\,\mathrm{d}V$
-$=\int_{\partial\Omega_{r}}|\nabla f|\langle\nabla\mathrm{R},\nabla f\rangle\mathrm{e}^{-cf}\,\mathrm{d}V$
+$=\int_{\partial\Omega_{r}}|\nabla f|\langle\nabla\mathrm{R},\nabla f\rangle\mathrm{e}^{-cf}\,\mathrm{d}S$
 $\quad+(c-1)\int_{\Omega_{r}}\mathrm{R}\langle\nabla\mathrm{R},\nabla f\rangle\mathrm{e}^{-cf}\,\mathrm{d}V+\frac{1}{2}\int_{\Omega_{r}}|\nabla\mathrm{R}|^{2}\mathrm{e}^{-cf}\,\mathrm{d}V$
 $=-\mathrm{e}^{-cr}\Big(\frac{c-1}{c}+1\Big)\int_{\Omega_{r}}|\nabla\mathrm{R}|^{2}\,\mathrm{d}V+\Big(\frac{c-1}{c}+\frac{1}{2}\Big)\int_{\Omega_{r}}|\nabla\mathrm{R}|^{2}\mathrm{e}^{-cf}\,\mathrm{d}V.$
-Therefore, we have:
+Therefore:
 $\int_{\Omega_{r}}$ $\mathrm{V}(\nabla f,\nabla f)\,\mathrm{e}^{-cf}\,\mathrm{d}V$
-$=\int_{\Omega_{r}}\nabla^{2}\mathrm{R}(\nabla f,\nabla f)\,\mathrm{e}^{-cf}\,\mathrm{d}V+\frac{1}{2}\int_{\Omega_{r}}\mathrm{R}\langle\nabla\mathrm{R},\nabla f\rangle\,\mathrm{e}^{-cf}\,\mathrm{d}V$
-$\quad-\frac{1}{4}\int_{\Omega_{r}}\mathrm{R}^{2}|\nabla f|^{2}\,\mathrm{e}^{-cf}\,\mathrm{d}V$
-$=\mathrm{e}^{-cr}\Big(\frac{c-1}{c}+1-\frac{1}{2c}\Big)\int_{\Omega_{r}}|\nabla\mathrm{R}|^{2}\,\mathrm{d}V-\Big(\frac{c-1}{c}+\frac{1}{2}-\frac{1}{2c}\Big)\int_{\Omega_{r}}|\nabla\mathrm{R}|^{2}\,\mathrm{e}^{-cf}\,\mathrm{d}V$
-$\quad-\frac{1}{4}\int_{\Omega_{r}}\mathrm{R}^{2}|\nabla f|^{2}\,\mathrm{e}^{-cf}\,\mathrm{d}V$
+$=\int_{\Omega_{r}}\nabla^{2}\mathrm{R}(\nabla f,\nabla f)\,\mathrm{e}^{-cf}\,\mathrm{d}V+\frac{1}{2}\int_{\Omega_{r}}\mathrm{R}\langle\nabla\mathrm{R},\nabla f\rangle\,\mathrm{e}^{-cf}\,\mathrm{d}V-\frac{1}{4}\int_{\Omega_{r}}\mathrm{R}^{2}|\nabla f|^{2}\,\mathrm{e}^{-cf}\,\mathrm{d}V$
 $=-\int_{\Omega_{r}}\Big(\frac{1}{4}\mathrm{R}^{2}|\nabla f|^{2}+\frac{3(c-1)}{2c}|\nabla\mathrm{R}|^{2}\Big)\,\mathrm{e}^{-cf}\,\mathrm{d}V+\frac{(4c-3)\mathrm{e}^{-cr}}{2c}\int_{\Omega_{r}}|\nabla\mathrm{R}|^{2}\,\mathrm{d}V.$
-Now taking the limit $r\rightarrow\infty$:
-$\int_{M}\mathrm{V}(\nabla f,\nabla f)\,\mathrm{e}^{-cf}\,\mathrm{d}V=-\int_{M}\Big(\frac{1}{4}\mathrm{R}^{2}|\nabla f|^{2}+\frac{3(c-1)}{2c}|\nabla\mathrm{R}|^{2}\Big)\,\mathrm{e}^{-cf}\,\mathrm{d}V.$
-8 Proof of Theorem 1.1(ii)
-Assume $\mathrm{U}_{ij}\equiv 0$.
-By Lemma 2.1, $\mathrm{tr}_{g}(\mathrm{U})=-\Delta\mathrm{R}$, hence $\Delta R\equiv 0$.
-Fix a regular value $r\in(0,1)$ of $f$ and write $\Omega_{r}=\{x\in M:f(x)\leq r\}$.
-For any $c>0$, Lemma 7.3 gives
-$0=\int_{\Omega_{r}}\mathrm{U}(\nabla f,\nabla f)\,\,\mathrm{d}V_{c}=\frac{1}{4}\int_{\Omega_{r}}\!\Big((\mathrm{R}^{2}-\mathrm{R})|\nabla f|^{2}+\frac{1}{c}|\nabla\mathrm{R}|^{2}\Big)\,\mathrm{d}V_{c}-\frac{\mathrm{e}^{-cr}}{4c}\int_{\Omega_{r}}\!|\nabla\mathrm{R}|^{2}\,\,\mathrm{d}V.$
-Multiplying by $4c$ and rearranging yields the basic identity
-$c\int_{\Omega_{r}}(\mathrm{R}^{2}-\mathrm{R})|\nabla f|^{2}\,\mathrm{e}^{-cf}\,\,\mathrm{d}V=\int_{\Omega_{r}}|\nabla\mathrm{R}|^{2}\,\big(\mathrm{e}^{-cr}-\mathrm{e}^{-cf}\big)\,\,\mathrm{d}V.$
-On the fixed bounded set $\Omega_{r}$ we have $0\leq f\leq r<1$, so all integrands are smooth and dominated by an $L^{1}$ function independent of $c$.
-Hence we may differentiate under the integral sign.
-Differentiating once gives
-$\int_{\Omega_{r}}(\mathrm{R}^{2}-\mathrm{R})|\nabla f|^{2}\,\mathrm{e}^{-cf}\,\,\mathrm{d}V-c\int_{\Omega_{r}}f\,(\mathrm{R}^{2}-\mathrm{R})|\nabla f|^{2}\,\mathrm{e}^{-cf}\,\,\mathrm{d}V=\int_{\Omega_{r}}\big(f\,\mathrm{e}^{-cf}-r\,\mathrm{e}^{-cr}\big)\,|\nabla\mathrm{R}|^{2}\,\,\mathrm{d}V.$
-Differentiating again yields
-$\int_{\Omega_{r}}\big(cf^{2}-2f\big)(\mathrm{R}^{2}-\mathrm{R})|\nabla f|^{2}\,\mathrm{e}^{-cf}\,\,\mathrm{d}V=\int_{\Omega_{r}}\big(r^{2}\,\mathrm{e}^{-cr}-f^{2}\,\mathrm{e}^{-cf}\big)\,|\nabla\mathrm{R}|^{2}\,\,\mathrm{d}V.$
-An easy induction shows that for each integer $n\geq 0$,
-$\int_{\Omega_{r}}\big(nf^{\,n-1}-cf^{\,n}\big)\,(\mathrm{R}^{2}-\mathrm{R})|\nabla f|^{2}\,\mathrm{e}^{-cf}\,\,\mathrm{d}V=\int_{\Omega_{r}}\big(r^{\,n}\,\mathrm{e}^{-cr}-f^{\,n}\,\mathrm{e}^{-cf}\big)\,|\nabla\mathrm{R}|^{2}\,\,\mathrm{d}V.$ (8)
-Setting $c=0$ in (8) and summing over $n\geq 0$, we use the geometric series sum formulas:
-$\sum_{n=0}^{\infty}nf^{\,n-1}=\frac{1}{(1-f)^{2}},\qquad\sum_{n=0}^{\infty}f^{\,n}=\frac{1}{1-f},\qquad\sum_{n=0}^{\infty}r^{\,n}=\frac{1}{1-r}.$
-This leads to the identity
-$\int_{\Omega_{r}}\Bigg(\frac{(\mathrm{R}^{2}-\mathrm{R})|\nabla f|^{2}}{(1-f)^{2}}+\frac{f-r}{(1-f)(1-r)}\,|\nabla\mathrm{R}|^{2}\Bigg)\,\,\mathrm{d}V=0.$ (9)
-On $\Omega_{r}$ we have $0\leq\mathrm{R}\leq f<1$, hence $\mathrm{R}^{2}-\mathrm{R}\leq 0$ and $f-r\leq 0$, while (1f)>0(1-f)>0 and (1r)>0(1-r)>0.
-Therefore both terms in the integrand of (9) are non-positive.
-Since the integral vanishes, both terms vanish identically on $\Omega_{r}$:
-[…]
-In particular $\nabla\mathrm{R}\equiv 0$ on $\Omega_{r}$.
-Together with $\Delta\mathrm{R}\equiv 0$, unique continuation implies $\mathrm{R}$ is constant on all of $M$.
-If $\mathrm{R}=0$ at some point then $\mathrm{R}\equiv 0$ everywhere and the soliton is isometric to the Gaussian soliton on $\mathbb{R}^{4}$.
-Otherwise $\mathrm{R}>0$ and the soliton is Einstein.
-9 Proof of Theorem 1.2 Outside the Cone
-Assume $3\beta\neq\alpha$, then $\mathrm{tr}_{g}\mathfrak{B}_{ij}=(3\beta-\alpha)\Delta\mathrm{R}=0$, hence $\Delta\mathrm{R}\equiv 0$.
-For any $c>0$ and a fixed regular value $r\in(0,1)$ for $f$, let $\Omega_{r}=\{f\leq r\}.$ We express the Bach-like tensor as
-$\mathfrak{B}_{ij}=6\beta\,\mathrm{B}_{ij}+(\alpha-3\beta)\,\mathrm{U}_{ij}$
-One can easily check that for the volume element $\mathrm{d}V_{c}$, the integration of the Bach tensor yields:
-$\int_{\Omega_{r}}\mathrm{B}(\nabla f,\nabla f)\,\,\mathrm{d}V_{c}=-\frac{c}{2}\int_{\Omega_{r}}|\mathrm{D}_{ijk}|^{2}\,\mathrm{d}V_{c}=\mathrm{F}(c).$
-So differentiating past the integral symbol yields
-$\frac{d^{n}}{dc^{n}}\Big|_{c=0}\mathrm{F}(c)=-\frac{1}{2}\int_{\Omega_{r}}nf^{n-1}\,|\mathrm{D}_{ijk}|^{2}\,\mathrm{d}V.$
-We will need to multiply both sides of the Bach-like equation by $c$ to proceed.
-$\frac{d^{n}}{dc^{n}}\Big(c\,\mathrm{F}(c)\Big)=n\,\mathrm{F}^{(n-1)}(c)+c\,\mathrm{F}^{(n)}(c),$
-hence
-$\frac{d^{n}}{dc^{n}}\Big|_{c=0}=n\,\mathrm{F}^{(n-1)}(0)=-\frac{1}{2}\int_{\Omega_{r}}n(n-1)\,f^{n-2}\,|\mathrm{D}_{ijk}|^{2}\,\mathrm{d}V$
-Summing over $n$ on the right-hand side, we obtain
-$\int_{\Omega_{r}}\frac{|\mathrm{D}_{ijk}|^{2}}{(1-f)^{3}}\,\mathrm{d}V.$ (10)
-Now assuming $\mathfrak{B}_{ij}\equiv 0$, and multiplying both sides by $c$:
-$0=6\beta\,c\int_{\Omega_{r}}\mathrm{B}(\nabla f,\nabla f)\,\mathrm{d}V_{c}+(\alpha-3\beta)\,c\int_{\Omega_{r}}\mathrm{U}(\nabla f,\nabla f)\,\mathrm{d}V_{c},$
-hence
-$0=\frac{-6\beta c^{2}}{2}\int_{\Omega_{r}}|\mathrm{D}_{ijk}|^{2}\,\mathrm{d}V_{c}+(\alpha-3\beta)\,c\int_{\Omega_{r}}\mathrm{U}(\nabla f,\nabla f)\,\mathrm{d}V_{c}.$
-Now differentiating with respect to $c$, evaluating each derivative at $c=0$ and summing over $n$ we finally obtain
-$\boxed{-6\beta\int_{\Omega_{r}}\frac{|\mathrm{D}_{ijk}|^{2}}{(1-f)^{3}}\,\mathrm{d}V+\frac{1}{4}(\alpha-3\beta)\int_{\Omega_{r}}\Bigg(\frac{(\mathrm{R}^{2}-\mathrm{R})|\nabla f|^{2}}{(1-f)^{2}}+\frac{f-r}{(1-f)(1-r)}\,|\nabla\mathrm{R}|^{2}\Bigg)\,\,\mathrm{d}V=0.}$
-Now as $0<r<1$, both terms in the second integrand are negative as in (9). If $\beta>0$ and $\alpha-3\beta>0$, or if $\beta<0$ and $\alpha-3\beta<0$, then
-$|\mathrm{D}_{ijk}|^{2}\equiv 0,\quad\mathrm{R}^{2}-\mathrm{R}\equiv 0,\quad\mathrm{and}\quad|\nabla\mathrm{R}|^{2}\equiv 0\quad\mathrm{on}\,\,\Omega_{r}.$
-Hence, $M$ must be Einstein or isometric to the Gaussian shrinking soliton.
-These ranges for $\alpha$ and $\beta$ give us the desired result for the region outside the cone $\mathcal{C}$, hence for all $(\alpha,\beta)$ aside from $(0,0)$ and the Bach-line.
+Taking the limit $r\rightarrow\infty$ gives the stated global identity. ∎
+8 Proof of Theorem 1.2
+Proof.
+Assume that $\mathrm{U}_{ij}=0$.
+By Lemma 2.1, $\Delta\mathrm{R}=0$.
+Let $p$ be the point where $f$ attains its minimum value $f_{\min}$.
+Then $\nabla f(p)=0$, so by (2), $\mathrm{R}(p)=f_{\min}$.
+Using $\mathrm{U}=0$ at $p$ with $\Delta\mathrm{R}=0$:
+$\mathrm{Rc}(p)\big(1-\mathrm{R}(p)\big)+\tfrac{1}{4}\mathrm{R}(p)\big(\mathrm{R}(p)-1\big)g(p)=0,$
+which factors as
+$\big(1-\mathrm{R}(p)\big)\!\left(\mathrm{Rc}(p)-\tfrac{1}{4}\mathrm{R}(p)\,g(p)\right)=0.$
+If $\mathrm{R}(p)\neq 1$ then $\mathrm{Rc}(p)=\tfrac{1}{4}\mathrm{R}(p)g(p)$, which gives
+$|\mathrm{Rc}(p)|^{2}=\tfrac{1}{4}\mathrm{R}^{2}(p).$
+Since $\nabla f(p)=0$, Lemma 7.2 gives $2|\mathrm{Rc}(p)|^{2}=\mathrm{R}(p)$. Substituting:
+$\tfrac{1}{2}\mathrm{R}^{2}(p)=\mathrm{R}(p)\implies\mathrm{R}(p)\big(\mathrm{R}(p)-2\big)=0.$
+Hence $\mathrm{R}(p)\in\{0,2\}$. Together with the $\mathrm{R}(p)=1$ case, we have $f_{\min}=\mathrm{R}(p)\in\{0,1,2\}$.
+Case $f_{\min}=0$. We have $\mathrm{R}(p)=0$, and the manifold is isometric to the Gaussian soliton by [8].
+Case $f_{\min}=2$. Using the soliton equation,
+$\mathrm{R}^{2}-2|\mathrm{Rc}|^{2}=\mathrm{R}^{2}-2\left|\tfrac{1}{2}g-\nabla^{2}f\right|^{2}=\mathrm{R}^{2}-2\mathrm{R}+2-2|\nabla^{2}f|^{2}.$
+At the minimum $p$ with $f_{\min}=2$, we have $\Delta f(p)=2-\mathrm{R}(p)=0$. Since $p$ is a minimum, $\nabla^{2}f(p)\geq 0$, and $\Delta f(p)=0$ forces $\nabla^{2}f(p)=0$. Thus:
+$\mathrm{R}^{2}(p)-2\mathrm{R}(p)+2-2|\nabla^{2}f(p)|^{2}=4-4+2-0=2>0.$
+By continuity, there exist $r>2$ and $\epsilon>0$ such that $\mathrm{R}^{2}-2\mathrm{R}+2-2|\nabla^{2}f|^{2}\geq\epsilon>0$ on $\Omega_{r}$. Therefore, under $\mathrm{U}\equiv 0$,
+$\int_{\Omega_{r}}(\mathrm{R}^{2}-2|\mathrm{Rc}|^{2})|\nabla f|^{2}\,\mathrm{e}^{-f}\,\mathrm{d}V=0$
+implies $|\nabla f|^{2}\equiv 0$ on $\Omega_{r}$, hence $\mathrm{R}=f-|\nabla f|^{2}=2$ on $\Omega_{r}$.
+By unique continuation for harmonic functions, $\mathrm{R}=2$ on all of $M$, and the soliton is Einstein by [8].
+For the converse, if $M$ is Einstein then $\mathrm{B}\equiv 0$ (as $\mathrm{B}$ vanishes on Einstein manifolds [4]) and $\mathrm{V}\equiv 0$ by Theorem 1.1, so $\mathrm{U}=2\mathrm{B}-\tfrac{1}{3}\mathrm{V}\equiv 0$.
+If $M$ is Gaussian then all curvature vanishes and $\mathrm{U}\equiv 0$.
+Case $f_{\min}=1$.
+At a minimum point $p$ the $\mathrm{U}$-tensor vanishes independently of any global assumption, and $\mathrm{V}$ cannot vanish.
+Coupled with the sign issue for the $\mathrm{U}$-vanishing integral identity, this case remains open; see Conjecture 1.3.
+∎
 9 Variational Origins of U and V
@@
 The approach taken by R.
-Bach’s [bach:1921] was to consider the first variation of the quadratic Riemannian functional
+Bach [2] was to consider the first variation of the quadratic Riemannian functional
 $\mathrm{S}_{W}=\int|\mathrm{W}|^{2}\,\,\mathrm{d}V_{g}.$
-As can be seen in A. Besse [besse:1987], one has
+As seen in A. Besse [3], one has
 $\mathrm{B}\,\,\propto\,\,\mathrm{grad}\left(\int|\mathrm{W}|^{2}\,\mathrm{d}V_{g}\right),$
-so the Bach tensor is a critical point of a Riemannian functional whose energy depends on how much a metric deviates from being conformally flat.
+so the Bach tensor is the gradient of a Riemannian functional whose energy depends on how much a metric deviates from being conformally flat.
 In this section, we identify Bach-like tensors as the Euler–Lagrange gradients of natural quadratic curvature functionals.
-We will work in the standard transverse–traceless (TT) gauge for the second variation, imposing for a variation $h$
+We work in the standard transverse-traceless (TT) gauge for the second variation, imposing for a variation $h$:
 $\mathrm{(TT)\,gauge:}\quad\mathrm{Div}_{g}(h)=0\quad\mathrm{and}\quad\mathrm{tr}_{g}(h)=0.$
-It is worth recalling the Lichnerowicz Laplacian as it will appear in our second variation formulas. This operator $\Delta_{L}$ acts on symmetric 2-tensors via the rule
-$(\Delta_{L}\,h)_{ij}=(\nabla^{*}\nabla\,h)_{ij}+2\,\mathrm{R}_{ikjl}\,h_{kl}-\mathrm{R}_{ik}\,h_{kj}-\mathrm{R}_{jk}\,h_{ki}.$
-Here $\nabla^{*}$ denotes the divergence operator or formal adjoint of the covariant derivative $\nabla$, and $\nabla^{*}\nabla$ is the rough Laplacian.
-Note that sign conventions for the Laplacian vary, but we will take the convention that
-$\nabla^{*}\nabla\,h=-\nabla^{k}\nabla_{k}\,h.$
+Recall the Lichnerowicz Laplacian acting on symmetric $2$-tensors:
+$(\Delta_{L}\,h)_{ij}=(\nabla^{*}\nabla\,h)_{ij}+2\,\mathrm{R}_{ikjl}\,h_{kl}-\mathrm{R}_{ik}\,h_{kj}-\mathrm{R}_{jk}\,h_{ki},$
+where $\nabla^{*}\nabla\,h=-\nabla^{k}\nabla_{k}\,h$.
 Lemma 9.1.
@@
 $\mathcal{F}_{\alpha,\beta}(g):=-\,\alpha\int_{M}|\mathrm{W}_{g}|^{2}\,\mathrm{d}V_{g}\;+\;\frac{\alpha/3-\beta}{2}\int_{M}\mathrm{R}_{g}^{2}\,\mathrm{d}V_{g}.$
-Then the $L^{2}$–gradient satisfies
-$\mathrm{grad}\,\mathcal{F}_{\alpha,\beta}\;=\;\alpha\,\mathrm{U}\;+\;\beta\,\mathrm{V}.$
-In particular, the Euler–Lagrange equation for $\mathcal{F}_{\alpha,\beta}$ is equivalent to the vanishing of the Bach-like tensor $\mathfrak{B}_{ij}=\alpha\,\mathrm{U}_{ij}+\beta\,\mathrm{V}_{ij}=0$.
+Then the $L^{2}$-gradient satisfies $\mathrm{grad}\,\mathcal{F}_{\alpha,\beta}=\alpha\,\mathrm{U}+\beta\,\mathrm{V}$.
+In particular, the Euler–Lagrange equation for $\mathcal{F}_{\alpha,\beta}$ is $\mathfrak{B}_{ij}=\alpha\,\mathrm{U}_{ij}+\beta\,\mathrm{V}_{ij}=0$.
 Proof.
-In dimension four, one has
+In dimension four:
 $\delta\!\int|\mathrm{W}|^{2}\,\mathrm{d}V_{g}=-2\!\int\langle\mathrm{B},\delta g\rangle\,\mathrm{d}V_{g}\quad\mathrm{and}\quad\delta\!\int\mathrm{R}^{2}\,\mathrm{d}V_{g}=-2\!\int\langle\mathrm{V},\delta g\rangle\,\mathrm{d}V_{g}.$
-These standard variational identities for quadratic functionals are well-established [e.g., [besse:1987]]. Using $\mathrm{B}=\tfrac{1}{2}\,\mathrm{U}+\tfrac{1}{6}\,\mathrm{V}$,
-$\delta\mathcal{F}_{\alpha,\beta}=\int\Big\langle\alpha\!\left(\tfrac{1}{2}\mathrm{U}+\tfrac{1}{6}\mathrm{V}\right)\;-\;\tfrac{\alpha/3-\beta}{2}\cdot 2\mathrm{V},\,\delta g\Big\rangle\,\mathrm{d}V_{g}=\int\langle\alpha\mathrm{U}+\beta\mathrm{V},\,\delta g\rangle\,\mathrm{d}V_{g}.$
-Hence $\mathrm{grad}\,\mathcal{F}_{\alpha,\beta}=\alpha\,\mathrm{U}+\beta\,\mathrm{V}$. ∎
+Using $\mathrm{B}=\tfrac{1}{2}\,\mathrm{U}+\tfrac{1}{6}\,\mathrm{V}$:
+$\delta\mathcal{F}_{\alpha,\beta}=\int\Big\langle\alpha\!\left(\tfrac{1}{2}\mathrm{U}+\tfrac{1}{6}\mathrm{V}\right)\;-\;\tfrac{\alpha/3-\beta}{2}\cdot 2\mathrm{V},\,\delta g\Big\rangle\,\mathrm{d}V_{g}=\int\langle\alpha\mathrm{U}+\beta\mathrm{V},\,\delta g\rangle\,\mathrm{d}V_{g}.\qed$
 Corollary 9.2.
 $\mathrm{V}\;=\;-\tfrac{1}{2}\,\mathrm{grad}\!\left(\int_{M}\mathrm{R}^{2}\,\,\mathrm{d}V_{g}\right),\qquad\mathrm{U}\;=\;-\,\mathrm{grad}\!\left(\int_{M}\big(|\mathrm{W}|^{2}-\tfrac{1}{6}\mathrm{R}^{2}\big)\,\,\mathrm{d}V_{g}\right).$
-Proposition 10.3 (Second variation of $\int R^{2}$ in TT gauge).
-Let $(M^{4},g)$ be closed. For a variation $g_{t}=g+t\,h+o(t)$ with
-$\mathrm{div}h=0,\qquad\mathrm{tr}_{g}h=0\quad\text{\emph{(TT gauge)}},$
-the second variation of $F(g):=\int_{M}\mathrm{R}^{2}\,\,\mathrm{d}V_{g}$ is
+Proposition 9.3 (Second variation of $\int\mathrm{R}^{2}$ in TT gauge).
+Let $(M^{4},g)$ be closed. For a TT variation $g_{t}=g+th+o(t)$, the second variation of $F(g):=\int_{M}\mathrm{R}^{2}\,\,\mathrm{d}V_{g}$ is
 $\delta^{2}F[h,h]=\int_{M}\Big[-\,\mathrm{R}\,|\nabla h|^{2}-2\mathrm{R}\,\mathrm{R}_{ikjl}\,h_{ij}h_{kl}+2\mathrm{R}\,\mathrm{R}_{ik}\,h_{ij}h_{kj}+2\langle\mathrm{Rc},h\rangle^{2}+\Big(\frac{1}{2}\mathrm{R}^{2}-2\,\Delta\mathrm{R}\Big)|h|^{2}\Big]\,\,\mathrm{d}V.$
 Proof.
-Write the first variation as $\delta F=-2\int\langle\mathrm{V},h\rangle\,\mathrm{d}V_{g}$. Then by differentiating once more, and using $\mathrm{div}h=0$, $\mathrm{tr}_{g}h=0$, the Hessian is
+Write the first variation as $\delta F=-2\int\langle\mathrm{V},h\rangle\,\mathrm{d}V_{g}$. Differentiating once more and using TT gauge:
 $\delta^{2}F[h,h]=-2\int_{M}\langle\mathrm{V}^{\prime}(h),h\rangle\,\mathrm{d}V_{g}.$
-We have the following standard TT-gauge variations:
+Expanding using the standard TT-gauge variations
 $\delta\mathrm{R}$ $=-\langle\mathrm{Rc},h\rangle,$
-$\delta(\nabla_{i}\nabla_{j}\mathrm{R})$ $=\nabla_{i}\nabla_{j}(\delta\mathrm{R})-\tfrac{1}{2}(\nabla_{i}h_{j}{}^{k}+\nabla_{j}h_{i}{}^{k}-\nabla^{k}h_{ij})\,\nabla_{k}\mathrm{R},$
 $\delta(\Delta\mathrm{R})$ $=\Delta(\delta\mathrm{R})-\langle h,\nabla^{2}\mathrm{R}\rangle,$
-$\delta\mathrm{Rc}$ $=\tfrac{1}{2}(\nabla^{*}\nabla h)_{ij}+(\mathrm{Rm}\!*h)_{ij}-(\mathrm{Rc}\circ h+h\circ\mathrm{Rc})_{ij}.$
-Using these, the Hessian can be written as
-$\delta^{2}F[h,h]=\int_{M}\Big(-\mathrm{R}\,\langle\Delta_{L}h,h\rangle+2\langle\mathrm{Rc},h\rangle^{2}+(\frac{1}{2}\mathrm{R}^{2}-2\Delta\mathrm{R})|h|^{2}\Big)\,\,\mathrm{d}V_{g}.$
-We expand the Lichnerowicz laplacian as $\Delta_{L}h=\nabla^{*}\nabla h+2\mathrm{Rm}\!*h-\mathrm{Rc}\circ h-h\circ\mathrm{Rc}$. Integration by parts yields the result. ∎
+$\delta\mathrm{Rc}$ $=\tfrac{1}{2}(\nabla^{*}\nabla h)_{ij}+(\mathrm{Rm}\!*h)_{ij}-(\mathrm{Rc}\circ h+h\circ\mathrm{Rc})_{ij},$
+and integrating by parts yields the result. ∎
 Remark 9.4.
@@
 Proposition 9.5 (Second variation at Einstein backgrounds).
-On an Einstein $4$-manifold $(M,g_{E})$ with $\mathrm{Rc}=\lambda g_{E}$ and scalar curvature $R=4\lambda$, the second variation of $\mathcal{F}_{\alpha,\beta}(g)$ in the TT-gauge is
+On an Einstein $4$-manifold $(M,g_{E})$ with $\mathrm{Rc}=\lambda g_{E}$ and $\mathrm{R}=4\lambda$, the second variation of $\mathcal{F}_{\alpha,\beta}(g)$ in TT gauge is
 $\delta^{2}\mathcal{F}_{\alpha,\beta}[h,h]=\int_{M}\!\Big\langle\big(\alpha\,\Delta_{L}^{2}+\frac{1}{2}\mathrm{R}\,(\beta-\tfrac{4}{3}\alpha)\,\Delta_{L}+\mathrm{R}^{2}(\tfrac{5}{36}\alpha-\tfrac{1}{4}\beta)\big)h,\;h\Big\rangle\,\mathrm{d}V.$
@@
 Proof.
-In TT gauge at an Einstein metric, the standard linearizations give
+In TT gauge at an Einstein metric:
 $\mathrm{B}^{\prime}(h)=\tfrac{1}{2}\big(\Delta_{L}-\tfrac{\mathrm{R}}{3}\big)\big(\Delta_{L}-\tfrac{\mathrm{R}}{6}\big)h,\qquad\mathrm{V}^{\prime}(h)=\frac{1}{2}\mathrm{R}\,\Delta_{L}h.$
-Since
-$\delta^{2}\!\int_{M}|\mathrm{W}|^{2}\,\mathrm{d}V_{g}=-2\int_{M}\!\langle\mathrm{B}^{\prime}(h),h\rangle\,\mathrm{d}V_{g}\quad\mathrm{and}\quad\delta^{2}\!\int_{M}\mathrm{R}^{2}\,\mathrm{d}V_{g}=-2\int_{M}\!\langle\mathrm{V}^{\prime}(h),h\rangle\,\mathrm{d}V_{g},$
 The claimed polynomial follows by expanding
-$\delta^{2}\mathcal{F}_{\alpha,\beta}[h,h]=-2\alpha\!\int_{M}\langle\mathrm{B}^{\prime}(h),h\rangle\,\mathrm{d}V_{g}\;-\;(\alpha/3-\beta)\!\int_{M}\langle\mathrm{V}^{\prime}(h),h\rangle\,\mathrm{d}V_{g}.$
+$\delta^{2}\mathcal{F}_{\alpha,\beta}[h,h]=-2\alpha\!\int_{M}\langle\mathrm{B}^{\prime}(h),h\rangle\,\mathrm{d}V_{g}\;-\;(\alpha/3-\beta)\!\int_{M}\langle\mathrm{V}^{\prime}(h),h\rangle\,\mathrm{d}V_{g}.\qed$
 Corollary 9.6 (Bach line and factorization).
-On the Bach line $\beta=\alpha/3$, one has
-$\delta^{2}\mathcal{F}_{\alpha,\beta}[h,h]=\alpha\int_{M}\left\langle\big(\Delta_{L}-\tfrac{\mathrm{R}}{3}\big)\big(\Delta_{L}-\tfrac{\mathrm{R}}{6}\big)h,\;h\right\rangle\,\mathrm{d}V_{g},$
-i.e. the Hessian factorizes into two shifted Lichnerowicz factors.
+On the Bach line $\beta=\alpha/3$:
+$\delta^{2}\mathcal{F}_{\alpha,\beta}[h,h]=\alpha\int_{M}\left\langle\big(\Delta_{L}-\tfrac{\mathrm{R}}{3}\big)\big(\Delta_{L}-\tfrac{\mathrm{R}}{6}\big)h,\;h\right\rangle\,\mathrm{d}V_{g}.$
 Corollary 9.7 (Local rigidity by spectral positivity).
-Let $(M^{4},g_{E})$ be Einstein with TT–spectrum of $\Delta_{L}$ contained in $\mu_{0},\infty)$.
-If $p_{\alpha,\beta}(\mu)>0$ for every TT–eigenvalue $\mu$ (equivalently, $\inf_{\mu\geq\mu_{0}}p_{\alpha,\beta}(\mu)>0$), then $g_{E}$ is a strict local minimizer of $\mathcal{F}_{\alpha,\beta}$ modulo diffeomorphisms and scaling.
-In particular, there exists a neighborhood of $g_{E}$ in which the equation $\alpha U+\beta V=0$ has no nontrivial solutions other than metrics isometric to $g_{E}$.
+Let $(M^{4},g_{E})$ be Einstein with TT-spectrum of $\Delta_{L}$ contained in $\mu_{0},\infty)$.
+If $p_{\alpha,\beta}(\mu)>0$ for every TT-eigenvalue $\mu$, then $g_{E}$ is a strict local minimizer of $\mathcal{F}_{\alpha,\beta}$ modulo diffeomorphisms and scaling.
+In particular, there exists a neighborhood of $g_{E}$ in which $\alpha\,\mathrm{U}+\beta\,\mathrm{V}=0$ has no nontrivial solutions other than metrics isometric to $g_{E}$.
 Remark 9.8 (Flat/Gaussian background).
-If $R=0=W$ (Euclidean space, or a flat compact quotient), then $B^{\prime}(h)=\tfrac{1}{2}\Delta_{L}^{2}h$ and $V^{\prime}(h)\equiv 0$ on TT. Hence
-$\delta^{2}\mathcal{F}_{\alpha,\beta}[h,h]=\alpha\int_{M}\langle\Delta_{L}^{2}h,h\rangle\,\,\mathrm{d}V=\alpha\int_{M}|\Delta h|^{2}\,\,\mathrm{d}V\quad(\text{TT}),$
-which is strictly positive for $\alpha>0$ on compactly supported (or suitably decaying/weighted) variations.
+If $\mathrm{R}=0=\mathrm{W}$, then $\mathrm{B}^{\prime}(h)=\tfrac{1}{2}\Delta_{L}^{2}h$ and $\mathrm{V}^{\prime}(h)\equiv 0$ on TT. Hence
+$\delta^{2}\mathcal{F}_{\alpha,\beta}[h,h]=\alpha\int_{M}|\Delta h|^{2}\,\,\mathrm{d}V,$
+which is strictly positive for $\alpha>0$ on compactly supported variations.